### v2

# Large deviations for invariant measures of the stochastic Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

Abstract
We prove the validity of a small noise large deviation principle for the family of invariant measures $\{\mu_{\epsilon}\}_{\epsilon>0}$ associated to the one dimensional stochastic Allen-Cahn equation with inhomogeneous Dirichlet boundary conditions, perturbed by unbounded multiplicative noise. The novelty is that the system is not strongly dissipative, and the noise intensity can be of linear growth while the drift term is of polynomial growth. Using L. Simon’s convergence theorem, we show that the dynamics of the noiseless system converge in large time to a minimiser of the Ginzburg-Landau energy functional, which is unique due to the boundary condition. We obtain an estimate of the invariant measure on a bounded set in the Sobolev space $W^{k^{\star},p^{\star}}$, where $k^{\star}p^{\star}>1$, and $p^{\star}$ is large, which is compactly embedded in the continuous function space. As a consequence of the LDP result, we show that $\mu_{\epsilon}$ concentrates around the unique minimiser with such boundary conditions exponentially fast when $\epsilon$ is sufficiently small.
MSC2020 subject classifications: Primary 60H15, 60F10, 35R60; secondary 37A50, 37L55.
Keywords: large deviation principle; invariant measures; stochastic Allen-Cahn equation; unbounded multiplicative noise; inhomogeneous boundary conditions.
1 Introduction
In this paper, we study the following equation
$\begin{cases}\partial_{t}u_{\epsilon}(t,\xi)=\Delta u_{\epsilon}(t,\xi)+u_{\epsilon}(t,\xi)-u_{\epsilon}(t,\xi)^{3}+\epsilon^{\frac{1}{2}}g(t,\xi,u_{\epsilon}(t,\xi))\partial_{t}W(t,\xi),\xi\in(-L,L),\\ u_{\epsilon}(0,\xi)=x(\xi),\\ u_{\epsilon}(t,\xi)=\mathrm{sgn}(\xi),\,|\xi|=L,\;t\geq 0,\end{cases}$ (1.1)
where $W(t,\xi):=\sum_{k\in\mathbb{N}}e_{k}(\xi)\beta_{k}(t)$, $\{e_{k}\}$ is a complete orthogonal basis in $H:=L^{2}(-L,L)$, and $\{\beta_{k}(t)\}$ is a sequence of independent real valued Brownian motions on the probability space $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P})$. Assume that the mapping $g:[0,\infty)\times[-L,L]\times\mathbb{R}\rightarrow\mathbb{R}$, is measurable, and for almost all $t\geq 0$, $g(t,\cdot,\cdot):[-L,L]\times\mathbb{R}\rightarrow\mathcal{L}(\mathbb{R})$, is continuous. Moreover, for any $\theta,\vartheta\in\mathbb{R}$ and almost all $t\geq 0$,
$\sup_{\xi\in[-L,L]}|g(t,\xi,\theta)-g(t,\xi,\vartheta)|\leq\Psi(t)|\theta-\vartheta|,$
for some $\Psi\in L^{\infty}_{loc}[0,\infty)$.
To give a precise meaning to the solution of the above equation, we set
$\psi(\xi)=\xi/L,\,\xi\in[-L,L]$ (1.2)
and denote by $S(t)$ the heat semigroup on (L,L)(-L,L) with zero boundary conditions at the endpoints. We denote by $\mathcal{E}$ the set of all $u\in C[-L,L]$ with u(L)=1,u(L)=1u(-L)=-1,\,u(L)=1 and by $E$ the space of all $u\in C[-L,L]$ with u(L)=u(L)=0u(-L)=u(L)=0. Note that $\mathcal{E}$ is not a linear space. Then, the mild solution of Eq.(1.1) is defined as $u_{\epsilon}(t):=\bar{u}_{\epsilon}(t)+\psi$, and $\bar{u}_{\epsilon}$ satisfies
$\bar{u}_{\epsilon}(t)$ $=S(t)(x-\psi)+\int_{0}^{t}S(t-s)\left[(\bar{u}_{\epsilon}(s)+\psi)-(\bar{u}_{\epsilon}(s)+\psi)^{3}\right]\,ds$
$\ \ \ \ +\epsilon^{\frac{1}{2}}\int_{0}^{t}S(t-s)G(s,\bar{u}_{\epsilon}(s))\,dW(s),$
is the mild solution of the following equation
$\begin{cases}\partial_{t}\bar{u}_{\epsilon}(t,\xi)=\Delta\bar{u}_{\epsilon}(t,\xi)+F(\bar{u}_{\epsilon}(t))(\xi)+\epsilon^{\frac{1}{2}}G(t,\bar{u}_{\epsilon}(t))\partial_{t}W(\xi),\xi\in(-L,L),\\ \bar{u}_{\epsilon}(0)=x-\psi,\\ \bar{u}_{\epsilon}(t,\xi)=0,\,|\xi|=L,\;t\geq 0.\end{cases}$ (1.3)
Here, $G(t,z)\phi(\xi):=g(t,\xi,z(\xi)+\psi(\xi))\phi(\xi)$ and $F(z)(\xi):=-(z(\xi)+\psi(\xi))^{3}+(z(\xi)+\psi(\xi))$, $z\in E$. Note that for each $z$ that satisfies the zero boundary condition, $F(z)$ also satisfies the zero boundary condition, so the mild solution of Eq.(1.3) is well defined.
According to [cerrai2003stochastic] (see Theorem 5.3, Theorem 5.5, and Proposition 6.1), for any initial datum $x\in E$ and $T>0$ , there exists a unique mild solution $\bar{u}_{\epsilon}^{x}\in L^{p}(\Omega;C([0,T];E))$ with $p>1$. Moreover, there exists $c_{p}>0$ such that
$\mathbb{E}\sup_{t\geq 0}\|\bar{u}^{x}_{\epsilon}(t)\|^{p}_{E}\leq c_{p}(1+\|x-\psi\|^{p}_{E}).$
Furthermore, due to the Krylov–Bogoliubov theorem, there exists a sequence $\{t_{n}\}\uparrow+\infty$ such that the sequence of probability measures
$\bar{\mu}_{\epsilon,n}(B):=\frac{1}{t_{n}}\int_{0}^{t_{n}}\mathbb{P}(\bar{u}_{\epsilon}^{0}(s)\in B)\,ds,\,B\in\mathcal{B}(E)$
converges weakly to some measure $\bar{\mu}_{\epsilon}$, which is an invariant measure of Eq.(1.3) (see [cerrai2003stochastic], Theorem 6.2). We assume the following condition,
Assumption 1.
Assume that
$\inf_{(t,\xi,\theta)\in[0,+\infty)\times[-L,L]\times\mathbb{R}}|g(t,\xi,\theta)|=g_{0}>0.$ (1.4)
By Theorem 3.1 of [maslowski2000probabilistic] and Chapter 7 of [da1996ergodicity] (see also Section 5 of [goldys200512]), under the above assumption, for every $\epsilon>0$ the invariant measure $\bar{\mu}_{\epsilon}$ is unique. We define an invariant measure $\mu_{\epsilon}$ for Eq.(1.1) by
$\mu_{\epsilon}(B+\psi):=\bar{\mu}_{\epsilon}(B),\;B\in\mathcal{B}(E).$ (1.5)
Here, $B+\psi:=\{x\in\mathcal{E}:x=y+\psi,y\in E\}$. The support of this measure is contained in $\mathcal{E}$. The purpose of this paper is to study the LDP problem for $\mu_{\epsilon}$.
The study of small noise LDP for invariant measures of SPDEs began in the work of Sowers [sowers1992large]. Cerrai and Röckner considered the stochastic reaction-diffusion equation with a strongly dissipative nonlinear term in a multi-dimensional space ([cerrai2005large]). Brzeźniak and Cerrai studied the stochastic Navier-Stokes equation on $\mathbb{T}^{2}$ with additive noise ([brzezniak2017large]). Wang investgated the stochastic reaction-diffusion equation on an unbounded domain under strongly dissipative conditions ([wang2024large]). Bai, Feng and Zhao studied the stochastic Burgers equation on the $(0,1)$ domain ([BAI2026111284]). In all these cases, the systems have 0 as their global attractor in the underlying space, following a dissipative argument. In this paper, the dynamics are only weakly dissipative and the existing arguments fail to work for our situation. The specific potential $V(u)=\frac{1}{4}(u^{2}-1)^{2}$ we consider in this paper has double wells with 2 critical points $\pm 1$ where the potential attains its minimum. Due to the boundary condition of our problem, there exists a unique solution for the following equation
$\begin{cases}\Delta m_{L}(\xi)=V^{\prime}(m_{L}(\xi)),\\ m_{L}(\pm L)=\pm 1.\end{cases}$ (1.6)
Moreover, the solution of the noiseless system $u_{0}(t)$ (i.e. $\epsilon=0$ in equation (1.1)) is the $L^{2}$ gradient flow of the energy functional
$\mathbf{E}_{L}(u):=\int_{-L}^{L}\left[\frac{|u^{\prime}(\xi)|^{2}}{2}+V(u(\xi))\right]\;d\xi,\;u(\pm L)=\pm 1.$ (1.7)
By using L. Simon’s convergence theorem, we show that for any $x\in E$, $u_{0}(t)$ converges to the minimizer of $\mathbf{E}_{L}$, which is the unique solution of (1.6), as time goes to infinity. The minimizer $m_{L}$ is the stationary profile with two pure phases of alloys, magnetization, or fluids coexisting at the two ends of a bounded interval. In these contexts, the boundary models the effects of two different materials, magnetization, or fluids applied to the endpoints.
Note that in the case where the noise is space-time white noise, the invariant measure $\mu$ has an exponential density with respect to the Brownian bridge measure [reznikoff2005invariant]. By the exponential tilting argument, as an alternative version of the Varadhan Lemma, the LDP for the invariant measure will follow immediately (see Theorem III.17 of [hollander2000large]). In the case where $L(\epsilon)$ depends on $\epsilon$ and goes to infinity as $\epsilon\rightarrow 0$, the limiting behaviour of the invariant measure $\mu_{\epsilon}$ depends on the rate at which $L(\epsilon)$ diverges. Such problems are considered by Weber [weber2010sharp], Otto-Weber-Westdickenburg [otto2014invariant], Bertini-Stella-Buttà [bertini2008dobrushin], and Bertini-Buttà-Gesù [bertini2025asymptotics]. For all the works mentioned above, the noise must be additive space-time white noise. In the case of the multiplicative noise we consider here, the density and tilted LDP arguments break down. As far as we know, this is the first paper to consider the LDP for the invariant measure of semilinear parabolic equations with a double-well potential and multiplicative noise.
The typical approach to proving the LDP for invariant measures of SPDEs with multiplicative noise involves a crucial step to obtain an exponential estimate for the invariant measure $\bar{\mu}$ outside of the compact subset $B_{W^{k^{\star},p^{\star}}(-L,L)}(0,R)$ of the space of continuous functions $E$. The noise we consider is multiplicative space-time white noise. The intensity of the noise only needs to be Lipschitz, so it is of linear growth. Note that we impose a nondegeneracy condition on the noise in order to analyze the relationship between the energy functional and the rate functional. It is worth mentioning that in Cerrai-Röckner [cerrai2005large], when the nonlinear term is cubic but still dissipative, the noise intensity cannot have linear growth. The exponential estimate for $\bar{\mu}$ is actually quite challenging in our case. In [cerrai2022large], the authors chose the compact subset to be a bounded set of the Sobolev space $H^{1}$, while Wang [wang2024large] dealt with trace-class additive noise in $H^{1}(\mathbb{R})$, and chose the compact subset as the intersection of bounded sets in $H^{1}$ and a weighted $L^{2}$ space. This was followed by the desired exponential estimate when applying Ito’s formula on $H^{1}$. However, this approach does not work in our case because of the non-dissipative nature of our system and the lower regularity and growth rate of the intensity of the noise.
Motivated by some ideas from the previous work [BAI2026111284], and using the smoothing property of the heat semigroup, we transform the estimate for the $W^{k^{\star},p^{\star}}$ norm of the solution of the SPDE (1.3) to that of the following stochastic convolution:
$\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t):=\int_{0}^{t}S(t-s)e^{-\lambda(t-s)}G(s,\bar{u}_{\epsilon}^{0}(s))\,dW(s).$ (1.8)
We choose to estimate the $W^{k^{\star},p^{\star}}$ norm of $\bar{u}^{0}_{\epsilon}$ for large $p^{\star}$ such that $p^{\star}k^{\star}>1$ to ensure that the Sobolev space $W^{k^{\star},p^{\star}}$ is compactly embedded in the space $E$. Then, the $W^{k^{\star},p^{\star}}$ estimate of $\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})$ can be obtained by the factorisation method under the condition that the noise intensity is bounded. When $g$ is unbounded, the estimate is split into two terms: probability conditioned on the set $\bar{u}^{0}_{\epsilon}(t)$ being bounded and unbounded under the $E$ norm, respectively. Conditioned on $\bar{u}^{0}_{\epsilon}(t)$ being bounded, the estimate is pinned to the case when $g$ is bounded from above. Then we use the LDP technique to obtain the exponential estimate for the conditional probability. The probability of $\bar{u}^{0}_{\epsilon}(t)$ being outside a bounded subset of $E$ is obtained from the variational principle, which was also used in Budhiraja-Dupuis-Maroulas [budhiraja2008large]. The advantage of this approach is that we avoid estimating the exponential expectation of $\bar{u}_{\epsilon}^{0}$ directly, which allows the noise intensity to grow linearly. We believe that this approach can be used to study large deviations for invariant measures of other SPDEs. With the exponential tightness result of $\mu_{\epsilon}$ and the fact that the corresponding noiseless dynamics has a unique global attractor, we are able to prove the LDP for $\mu_{\epsilon}$. Moreover, as a corollary, it is shown that $\mu_{\epsilon}$ concentrates around $m_{L}$, which is the non-trivial minimiser of $\mathbf{E}_{L}$, exponentially fast in $\epsilon$.
2 LDP for dynamics
The large deviation result for the dynamics of Eq.(1.3) has been well studied by Cerrai and Röckner in [cerrai2004large]. We formulate their results (see [cerrai2004large], Theorems 6.2 and 6.3) to obtain the Freidlin-Wentzell uniform large deviations principle (FWULDP) for the dynamics of Eq.(1.3):
1. For any $z\in C([0,T];E)$ and $\delta,\gamma>0$, there exists $\epsilon_{0}>0$ such that for any $\epsilon<\epsilon_{0}$,
$\mathbb{P}(\|\bar{u}_{\epsilon}^{x}-z\|_{C([0,T];E)}<\delta)\geq\exp\left(-\frac{I^{x}_{T}(z)+\gamma}{\epsilon}\right),$ (2.1)
uniformly with respect to $\|x\|_{E}\leq R$.
2. For any $r\geq 0$ and $\delta,\gamma>0$, there exists $\epsilon_{0}>0$ such that for any $\epsilon<\epsilon_{0}$,
$\mathbb{P}(dist_{C([0,T];E)}(\bar{u}_{\epsilon}^{x},K_{T}^{x}(r))\geq\delta)\leq\exp\left(-\frac{r-\gamma}{\epsilon}\right),$ (2.2)
uniformly with respect to $\|x\|_{E}\leq R$, where $K_{T}^{x}(r):=\{z\in C([0,T];E):I_{T}^{x}(z)\leq r\}$.
For every $z\in C({0,T};E)$, the rate functional $I_{T}^{x}$ is given by
$I_{T}^{x}(z):=\frac{1}{2}\inf\left\{\int_{0}^{T}\|f(t)\|_{H}^{2}\,dt;f\in L^{2}(0,T;H),z=z^{x}_{f}\right\},$
with the convention that $\inf\varnothing=+\infty$. Here, $z^{x}_{f}$ is the solution of the following skeleton equation,
$\begin{cases}\partial_{t}z(t,\xi)=\Delta z(t,\xi)+F(z(t))(\xi)+G(t,z(t,\xi))f(t)(\xi),\xi\in(-L,L),\\ z(0,\xi)=x(\xi),\\ z(t,\xi)=0,\;|\xi|=L,\;t\geq 0,\end{cases}$ (2.3)
for a given function $f\in L^{2}(0,T;H)$. Theorem 4.1 in [cerrai2004large] implies that for any $f\in L^{2}(0,T;H)$ and $x\in E$, there exists a unique mild solution to Eq.(2.3) in $C([0,T];E)$. Thus, $I_{T}^{x}$ is well defined. Moreover, according to Theorem 5.1 of [cerrai2004large], the level sets $K_{T}^{x}(r)$ are compact. For any fixed $\phi\in L^{2}(0,T;H)$ and $z\in C([0,T];E)$ we define
$Z_{\phi}(z)(t):=\int_{0}^{t}S(t-s)G(s,z(s))\phi(s)\,ds,\;t\in[0,T].$ (2.4)
According to Section 4 of [cerrai2004large],
$\|Z_{\phi}(z)\|_{C([0,T];E)}\leq c(T)\left(1+\|z\|_{C([0,T];E)}\right)\|\phi\|_{L^{2}(0,T;H)},$ (2.5)
and
$\|Z_{\phi}(z_{1})-Z_{\phi}(z_{2})\|_{C([0,T];E)}\leq c(T)\left(\|z_{1}-z_{2}\|_{C([0,T];E)}\right)\|\phi\|_{L^{2}(0,T;H)},$ (2.6)
for some continuous increasing function $c(t)$ with $c(0)=0$. Then, there exists $t_{0}>0$ such that $c(t_{0})\|\phi\|_{L^{2}(0,T;H)}\leq\frac{1}{2}$. Since this argument can be repeated in the time intervals $nt_{0},(n+1)t_{0}],\;n\in\mathbb{N}$, we get
$\|Z_{\phi}(z_{1})-Z_{\phi}(z_{2})\|_{C([0,T];E)}\leq\frac{1}{2}\|z_{1}-z_{2}\|_{C([0,T];E)},$ (2.7)
in the whole time interval $[0,T]$.
Before proving the LDP for $\mu_{\epsilon}$, we also need to show the Dembo-Zeitouni uniform large deviations principle (DZULDP). The following definition can be found in [dembo2009large].
Definition 2.1.
Let $\mathcal{M}$ be a Banach space, and $D\subset\mathcal{M}$ be some non-empty set. Suppose that for each $x\in D$, $\{\mu^{x}_{\epsilon}\}_{\epsilon>0}$ is a family of probability measures on $\mathcal{M}$ and the action function $I^{x}:\mathcal{M}\rightarrow[0,+\infty]$ is a good rate function. The family $\{\mu^{x}_{\epsilon}\}_{\epsilon>0}$ is said to satisfy a Dembo-Zeitouni uniform large deviation principle (DZULDP) in $\mathcal{M}$, with rate function $I^{x}$, uniformly with respect to $x\in D$, if the following estimates hold:
(Lower bound) For any $\gamma>0$ and open set $G\subset\mathcal{M}$, there exists $\epsilon_{0}>0$ such that
$\inf_{x\in D}\mu^{x}_{\epsilon}(G)\geq\exp\left(-\frac{1}{\epsilon}\left[\sup_{y\in D}\inf_{u\in G}I^{y}(u)+\gamma\right]\right),$
for any $\epsilon\leq\epsilon_{0}$.
(Upper bound) For any $\gamma>0$ and closed set $F\subset\mathcal{M}$, there exists $\epsilon_{0}>0$ such that
$\sup_{x\in D}\mu^{x}_{\epsilon}(F)\leq\exp\left(-\frac{1}{\epsilon}\left[\inf_{y\in D}\inf_{u\in F}I^{y}(u)-\gamma\right]\right),$
for any $\epsilon\leq\epsilon_{0}$.
Corollary 2.2.
Let $D\subset E$ be a compact set. If $\bar{u}_{\epsilon}^{x}$ is the mild solution of Eq.(1.3), then the family $\left\{\mathcal{L}(\bar{u}_{\epsilon}^{x})\right\}_{\epsilon>0}$ satisfies a Dembo-Zeitouni uniform large deviations principle in $C([0,T];E)$ with good rate function $I_{T}^{x}$, uniformly with respect to $x\in D$.
Proof.
Since we have FWULDP, according to Theorem 2.7 of [salins2019equivalences], it suffices to show that for any $x_{n}\rightarrow x\in E$,
$\lim_{n\rightarrow\infty}\max\left(\sup_{z\in K_{T}^{x_{n}}(r)}d_{C([0,T];E)}(z,K_{T}^{x}(r)),\sup_{z\in K_{T}^{x}(r)}d_{C([0,T];E)}(z,K_{T}^{x_{n}}(r))\right)=0.$
This is implied by the continuity of the skeleton function with respect to the initial conditions. To see this, first since $F$ is locally Lipschitz continuous on $E$, there exists $C_{R}>0$ such that
$\|F(x)-F(y)\|_{E}\leq C_{R}\|x-y\|_{E},$
for any $x,y\in B_{R}(E)$. Moreover, according to Theorem 4.1 of [cerrai2004large], for $x,y\in E$ and $\phi,\varphi\in L^{2}(0,T;H)$, there exists $R$ such that
$\max\{\|z_{\phi}^{x}\|,\|z_{\varphi}^{y}\|\}\leq c(T,\|\phi\|_{L^{2}(0,T;H)}+\|\varphi\|_{L^{2}(0,T;H)})(1+\|x\|_{E}+\|y\|_{E})<R.$
We set $\rho=z^{x}_{\phi}-z^{y}_{\varphi}.$ Then, for any $t\in[0,T]$
$\|\rho(t)\|_{E}$ $\leq\|S(t)(x-y)\|_{E}+\int_{0}^{t}\|S(t-s)\left(F(z_{\phi}^{x})(s)-F(z_{\varphi}^{y})(s)\right)\|_{E}\,ds$
$\ \ \ \ +\|Z_{\phi}(z_{\phi}^{x})-Z_{\varphi}(z_{\varphi}^{y})\|_{E}$
$\leq\|x-y\|_{E}+\int_{0}^{t}C_{R}\|\rho(s)\|_{E}\,ds+\|Z_{\phi-\varphi}(z_{\varphi}^{y})(t)\|_{E}$
$\ \ \ \ +\|Z_{\phi}(z_{\phi}^{x})(t)-Z_{\phi}(z_{\varphi}^{y})(t)\|_{E}$
$\leq\|x-y\|_{E}+\int_{0}^{t}C_{R}\|\rho\|_{C([0,s];E)}\,ds+c(T)(R+1)\|\phi-\varphi\|_{L^{2}(0,T;H)}$
$\ \ \ \ +\frac{1}{2}\|\rho\|_{C([0,T];E)}.$
Here, the last inequality is due to (2.5) and (2.7). Thus,
$\|\rho\|_{C([0,T];E)}$ $\leq 2\|x-y\|_{E}+2\int_{0}^{T}C_{R}\|\rho\|_{C([0,t];E)}\,dt$
$\ \ \ \ +2c(T)(R+1)\|\phi-\varphi\|_{L^{2}(0,T;H)}.$
By the Grönwall inequality, this implies
$\|z_{\phi}^{x}-z_{\varphi}^{y}\|_{C([0,T];E)}\leq c(T,\|x\|_{E},\|\phi\|_{L^{2}(0,T;H)})\left(\|x-y\|_{E}+\|\phi-\varphi\|_{L^{2}(0,T;H)}\right).$ (2.8)
Therefore, for any $z\in K_{T}^{x}$, there exists $\phi$ such that $z=z_{\phi}^{x}$ and $\frac{1}{2}\|\phi\|_{L^{2}(0,T;H)}\leq r$. Then, we have
$d_{C([0,T];E)}(z,K_{T}^{x_{n}}(r))\leq d_{C([0,T];E)}(z_{\phi}^{x},z_{\phi}^{x_{n}})\leq c(T,\|x\|_{E},r)\|x-x_{n}\|_{E},$
which implies the result. ∎
3 The rate functional of LDP for invariant measures
3.1 Energy functional and quasi-potential
In this section, we study the rate functional $U_{L}$ for invariant measures. We first introduce the energy functional $\mathbf{E}_{L}$. For functions $u:[-L,L]\rightarrow\mathbb{R}$ with boundary conditions $u(\pm L)=\pm 1$, we consider the energy functional (1.7) with
$V(u)=\frac{1}{4}(u^{2}-1)^{2}.$ (3.1)
According to Proposition 3.1 of [bertini2008dobrushin], the functional $\mathbf{E}_{L}$ has a unique minimizer $m_{L}$ in $\mathcal{E}$ and $m_{L}$ is the unique solution for the equation
$\begin{cases}\Delta m(\xi)=m^{3}(\xi)-m(\xi),\xi\in(-L,L)\\ m(\xi)=\mathrm{sgn}(\xi),|\xi|=L.\end{cases}$ (3.2)
Equation (3.2) yields that, for $\xi\in(-L,L)$
$m^{\prime}(\xi)^{2}=\frac{1}{2}(m^{2}(\xi)-1)^{2}+e_{L},$ (3.3)
where $e_{L}>0$ depends on the spatial scale $L$. Taking (3.3) into account in manipulating (1.7), for any $u\in\mathcal{E}$, we have
$\begin{split}\mathbf{E}_{L}(u)&=\int_{-L}^{L}\left[\frac{|u^{\prime}(\xi)|^{2}}{2}+\frac{1}{4}(u^{2}(\xi)-1)^{2}\right]\;d\xi\\ &=\int_{-L}^{L}\Bigg[\frac{1}{2}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\\ &\ \ \ \ +\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\cdot u^{\prime}(\xi)-\frac{1}{2}e_{L}\Bigg]\;d\xi\\ &=\frac{1}{2}\int_{-L}^{L}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\;d\xi\\ &\ \ \ \ +\int_{-1}^{1}\sqrt{\frac{1}{2}(u^{2}-1)^{2}+e_{L}}\;du-Le_{L}.\end{split}$ (3.4)
Then,
$\mathbf{E}_{L}(m_{L})=\int_{-1}^{1}\sqrt{\frac{1}{2}(u^{2}-1)^{2}+e_{L}}\;du-Le_{L}$ (3.5)
and for any $u\in\mathcal{E}$,
$\mathbf{E}_{L}(u)\geq\mathbf{E}_{L}(m_{L}).$
We denote for any $\bar{u}\in E$
$\mathbf{E}_{L}^{*}(\bar{u})=\mathbf{E}_{L}(\bar{u}+\psi)-\mathbf{E}_{L}(m_{L})$ (3.6)
so that $\mathbf{E}_{L}^{*}(m_{L}-\psi)=0$. Note that $m_{L}-\psi$ is the only solution for the equation
$\begin{cases}\Delta z(\xi)+F(z)(\xi)=0,\\ z(-L)=z(L)=0.\end{cases}$ (3.7)
One can prove the following proposition which will be needed in this paper:
Proposition 3.1.
There exists $\eta>0$ such that if $\bar{u}\in E$ and $\mathbf{E}_{L}^{*}(\bar{u})\leq\eta$, then
$\|\bar{u}-(m_{L}-\psi)\|_{E}\leq C\sqrt{\eta}.$
Proof.
For $\bar{u}\in E$, we denote $u=\bar{u}+\psi$. According to (3.4), (3.5), and (3.6),
$\mathbf{E}_{L}^{*}(\bar{u})=\frac{1}{2}\int_{-L}^{L}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\;d\xi.$
Let us write
$u^{\prime}(\xi)=V(u(\xi))+\phi(\xi),u(-L)=-1,$
where $V(\cdot)=\sqrt{\frac{1}{2}((\cdot)^{2}-1)^{2}+e_{L}}$ is locally Lipschitz. Assume $\mathbf{E}_{L}^{*}(\bar{u})\leq\eta$, then $\phi$ satisfies $\int_{-L}^{L}\phi^{2}(\xi)\;d\xi\leq 2\eta$. By the Cauchy-Schwarz inequality,
$\int_{-L}^{L}|\phi(\xi)|\;d\xi\leq 2\sqrt{L\eta}.$
Moreover, from (3.4)
$\eta\geq\mathbf{E}_{L}^{*}(\bar{u})$ $\geq\int_{\inf_{\xi\in[-L,L]}\bar{u}(\xi)}^{\sup_{\xi\in[-L,L]}\bar{u}(\xi)}V(u)\;du-Le_{L}-\left(\int_{-1}^{1}V(u)\;du-Le_{L}\right)$
$\geq\int_{\inf_{\xi\in[-L,L]}\bar{u}(\xi)}^{-1}V(u)\;du+\int_{1}^{\sup_{\xi\in[-L,L]}\bar{u}(\xi)}V(u)\;du$
This implies that we can choose $\eta$ sufficiently small such that $\sup_{\xi\in[-L,L]}|u(\xi)|\leq 2$. Thus, by (3.3) and the locally Lipchitz property of $V(\cdot)$, one obtains
$\frac{d}{d\xi}|u(\xi)-m_{L}(\xi)|$ $\leq|V(u(\xi))-V(m_{L}(\xi))|+|\phi(\xi)|$
$\leq C|u(\xi)-m_{L}(\xi)|+|\phi(\xi)|.$
By Grönwall’s inequality, one can choose $\eta$ small enough such that
$\|\bar{u}-(m_{L}-\psi)\|_{E}=\sup_{\xi\in[-L,L]}|u(\xi)-m_{L}(\xi)|\leq 2\sqrt{L\eta}e^{2LC}.$
This completes the proof. ∎
Now, we can define the rate functional. The quasi-potential $U_{L}:E\rightarrow\mathbb{R}$ for LDP of $\bar{\mu}_{\epsilon}$ is defined as
$U_{L}(\zeta):=\inf\left\{I_{T}^{m_{L}-\psi}(z):T>0,z\in C([0,T];E),z(0)=m_{L}-\psi,z(T)=\zeta\right\}.$ (3.8)
Note that the level sets of $\mathbf{E}_{L}^{*}$ are compact (see Proposition 3.1 of [bertini2008dobrushin]). By the Sobolev embedding theorem, $\mathbf{E}_{L}^{*}(\bar{u})$ is finite if and only if $\bar{u}\in H^{1}(-L,L)$. For $\bar{u},h\in H^{1}(-L,L)$, denote by $D_{h}\mathbf{E}_{L}^{*}(\bar{u})$ the Fréchet derivative of $\mathbf{E}_{L}^{*}$ on $\bar{u}$, then
$D_{h}\mathbf{E}_{L}^{*}(\bar{u})$ $=\lim_{t\rightarrow 0}\frac{\mathbf{E}_{L}^{*}(\bar{u}+th)-\mathbf{E}_{L}^{*}(\bar{u})}{t}$
$=\int_{-L}^{L}[-(\bar{u}+\psi)(\xi)+(\bar{u}+\psi)^{3}(\xi)]h(\xi)\;d\xi$
$\ \ \ \ +\int_{-L}^{L}(\bar{u}+\psi)^{\prime}(\xi)h^{\prime}(\xi)\;d\xi$
$=-\int_{-L}^{L}[\Delta\bar{u}(\xi)+(\bar{u}+\psi)(\xi)-(\bar{u}+\psi)^{3}(\xi)]h(\xi)\;d\xi$
Here, the last equality is due to the integration by parts and the fact that $\Delta\psi=0$. This implies that
$D_{h}\mathbf{E}_{L}^{*}(u)=\langle D\mathbf{E}_{L}^{*}(u),h\rangle_{H^{-1},H^{1}},$
where
$D\mathbf{E}_{L}^{*}(u)=-[\Delta u+(u+\psi)-(u+\psi)^{3}].$ (3.9)
The functions $U_{L}$ and $\mathbf{E}_{L}^{*}$ have the following connection.
Proposition 3.2.
For any $\zeta\in E$, there exists $C>0$ such that
$\mathbf{E}_{L}^{*}(\zeta)\leq CU_{L}(\zeta).$ (3.10)
Proof.
Suppose $U_{L}(\zeta)=+\infty$, then our claim is clearly true. Therefore, we assume $U_{L}(\zeta)<\infty$. Then, there exist some $T>0$ and $z\in C([0,T];E)$ satisfying $z(0)=m_{L}-\psi$, $z(T)=\zeta$ such that $z$ is a solution of equation (2.3) with $f\in L^{2}(0,T;H)$ such that
$I_{T}^{m_{L}-\psi}(z)=\frac{1}{2}\|f\|_{L^{2}(0,T;H)}^{2}<+\infty.$
Multiplying both parts of equation (2.3) by $D\mathbf{E}^{*}_{L}(z(t))$ and integrating with respect to $\xi$ in [L,L][-L,L], by the chain rule, one has
$\frac{d\;\mathbf{E}^{*}_{L}(z(t))}{dt}$ $=-\int_{-L}^{L}|\Delta z(t,\xi)+F(z(t))(\xi)|^{2}\;d\xi$
$\ \ \ \ -\int_{-L}^{L}G(t,z(t))f(t,\xi)\cdot[\Delta z(t,\xi)+F(z(t))(\xi)]\;d\xi$
$\leq-\int_{-L}^{L}|\Delta z(t,\xi)+F(z(t))(\xi)|^{2}\;d\xi+\int_{-L}^{L}|\Delta z(t,\xi)+F(z(t))(\xi)|^{2}\;d\xi$
$\ \ \ \ +\frac{1}{4}C(\|z(t)\|_{E}+1)^{2}\|f(t)\|_{H}^{2}.$
Integrating by $t$, one gets
$\mathbf{E}_{L}^{*}(\zeta)-\mathbf{E}^{*}_{L}(z(0))\leq\frac{C}{2}(\|z\|_{C([0,T];E)}^{2}+1)I_{T}(z).$ (3.11)
Following the same procedure as in the proof of Proposition 6.1 of [cerrai2003stochastic], we have
$\sup_{t\geq 0}\|z(t)\|_{E}\leq C(\|f\|_{L^{2}([0,+\infty);H)}+1).$ (3.12)
Since $\mathbf{E}_{L}^{*}(m_{L}-\psi)=0$, taking the infimum over $T$ and $z\in C([0,T];E)$,
$\mathbf{E}_{L}^{*}(\zeta)\leq CU_{L}(\zeta).$
The above proposition shows that the energy functional is controlled by the rate functional. A special case for the skeleton equation is $f=0$. The following lemma states the global dynamics of $z_{0}^{x}$.
Lemma 3.3.
For any $x\in H$, let $z_{0}^{x}$ be the solution of skeleton equation (2.3) with $f=0$. Then,
$\lim_{t\rightarrow+\infty}\|z_{0}^{x}(t)-(m_{L}-\psi)\|_{H^{1}(-L,L)}=0,$ (3.13)
and
$\lim_{t\rightarrow+\infty}\mathbf{E}_{L}^{*}(z_{0}^{x}(t))=0.$ (3.14)
Proof.
Multiplying both sides of equation (2.3) by $z_{0}^{x}(t)$ and integrating with respect to $\xi$, we have
$\ \ \ \ \frac{1}{2}\partial_{t}\|z_{0}^{x}(t)\|_{H}^{2}+\|z_{0}^{x}(t)\|_{H^{1}(-L,L)}^{2}$
$=\int_{-L}^{L}-(z_{0}^{x}(t,\xi)+\psi(\xi))^{3}\cdot z_{0}^{x}(t,\xi)+(z_{0}^{x}(t,\xi)+\psi(\xi))\cdot z_{0}^{x}(t,\xi)\;d\xi$
$\leq-\|z_{0}^{x}(t)\|_{H}^{2}+CL.$
By Grönwall’s inequality,
$\|z_{0}^{x}(t)\|_{H}^{2}+\int_{0}^{t}\|z_{0}^{x}(s)\|_{H^{1}(-L,L)}^{2}\;ds\leq C(t,L)(\|x\|_{H}^{2}+1).$
This implies that there exists $t^{\prime}>0$ such that $z_{0}^{x}(t^{\prime})\in H^{1}(-L,L)$. Multiplying both parts of equation (2.3) by $D\mathbf{E}^{*}_{L}(z^{x}_{0}(t))$ and integrating with respect to $\xi$ in [L,L][-L,L]. Since $f=0$, by the chain rule
$\frac{d\;\mathbf{E}^{*}_{L}(z_{0}^{x}(t))}{dt}=-\|\Delta z_{0}^{x}(t,\cdot)+F(z_{0}^{x}(t))(\cdot)\|_{H}^{2}.$
Thus, for any $t>t^{\prime}$,
$\mathbf{E}_{L}^{*}(z_{0}^{x}(t^{\prime}))-\mathbf{E}_{L}^{*}(z_{0}^{x}(t))=\int_{t^{\prime}}^{t}\|\Delta z_{0}^{x}(s,\cdot)+F(z_{0}^{x}(s))(\cdot)\|_{H}^{2}\;ds\geq 0.$ (3.15)
Since $\mathbf{E}_{L}^{*}(\zeta)\geq 0$, for any $\zeta\in E$, then for all $t>0$
$\int_{t^{\prime}}^{t}\|\Delta z_{0}^{x}(s,\cdot)+F(z_{0}^{x}(s))(\cdot)\|_{H}^{2}\;ds\leq\mathbf{E}_{L}^{*}(z_{0}^{x}(t^{\prime}))<+\infty,$ (3.16)
and
$\mathbf{E}_{L}^{*}(z_{0}^{x}(t^{\prime}))\geq\mathbf{E}^{*}_{L}(z^{x}_{0}(t))\geq\frac{1}{2}\|z_{0}^{x}(t)+\psi\|_{H^{1}}^{2}-\mathbf{E}_{L}(m_{L})\geq\|z_{0}^{x}(t)\|_{H^{1}}^{2}-\frac{1}{L}-\mathbf{E}_{L}(m_{L}).$
By compactness of the Sobolev embedding from $H^{1}(-L,L)$ to $E$ and (3.16), there exists a subsequences $z_{0}^{x}(t_{k})$ and $z_{L}\in E$ such that $z_{0}^{x}(t_{k})$ converges to $z_{L}$ weakly in $H^{1}(-L,L)$, strongly in $E$ as $t_{k}\rightarrow\infty$ and
$\lim_{k\rightarrow+\infty}\|\Delta z_{0}^{x}(t_{k})+F(z_{0}^{x}(t_{k}))\|_{H}=0.$
Since $F$ is locally Lipschitz, for $\tilde{h}\in H$
$\lim_{k\rightarrow+\infty}\langle F(z_{0}^{x}(t_{k}))-F(z_{L}),\tilde{h}\rangle_{H}\leq C\lim_{k\rightarrow+\infty}\langle|z_{0}^{x}(t_{k})-z_{L}|,|\tilde{h}|\rangle_{H}=0.$
Then, for $h\in H^{1}(-L,L)$,
$\lim_{k\rightarrow+\infty}\langle\Delta z_{0}^{x}(t_{k})+F(z_{0}^{x}(t_{k})),h\rangle_{H}$ $=\lim_{k\rightarrow+\infty}\left(\langle\nabla z_{0}^{x}(t_{k}),\nabla h\rangle_{H}+\langle F(z_{0}^{x}(t_{k})),h\rangle_{H}\right)$
$=\langle\nabla z_{L},\nabla h\rangle_{H}+\langle F(z_{L}),h\rangle_{H}$
$=\langle\Delta z_{L}+F(z_{L}),h\rangle=0.$
This implies
$\Delta z_{L}+F(z_{L})=0.$
Due to the uniqueness of the solution for equation (3.7), one obtains
$z_{L}=m_{L}-\psi.$
Moreover,
$\lim_{k\rightarrow+\infty}\|z_{0}^{x}(t_{k})-z_{L}\|_{H^{1}(-L,L)}^{2}$ $=\lim_{k\rightarrow+\infty}\langle\Delta z_{0}^{x}(t_{k})-\Delta z_{L},z_{0}^{x}(t_{k})-z_{L}\rangle_{H}$
$=\lim_{k\rightarrow+\infty}\langle\Delta z_{0}^{x}(t_{k})+F(z_{0}^{x}(t_{k})),z_{0}^{x}(t_{k})-z_{L}\rangle_{H}$
$\ \ \ \ +\lim_{k\rightarrow+\infty}\langle F(z_{L})-F(z_{0}^{x}(t_{k})),z_{0}^{x}(t_{k})-z_{L}\rangle_{H}$
$=0.$
Here, the Cauchy-Schwarz inequality and the local Lipschitz continuity of $F$ are used when the above two limits are taken. Therefore, $z_{0}^{x}(t_{k})$ converges to $m_{L}-\psi$ in $H^{1}(-L,L)$ and $\bigcup_{t\geq 1}\{z_{0}^{x}(t,\cdot)\}$ is precompact in $H^{1}(-L,L)$.
The above arguments show that the assumptions in Theorem 1.1 of [jendoubi1998simple] are satisfied. Applying L.Simon’s convergence theorem, we have that $z(t)$ is convergent to one equilibrium of the system which satisfies equation (3.7) as $t\rightarrow\infty$. By the uniqueness of limit,
$\lim_{t\rightarrow+\infty}\|z_{0}^{x}(t)-(m_{L}-\psi)\|_{H^{1}(-L,L)}=0.$
Thus, by the Sobolev embedding,
$\lim_{t\rightarrow+\infty}\mathbf{E}_{L}^{*}(z_{0}^{x}(t))=\mathbf{E}_{L}^{*}(m_{L}-\psi)=0.$
Proposition 3.4.
Under Assumption 1, for any $\zeta\in E$,
$g_{0}\cdot U_{L}\leq 2\mathbf{E}^{*}_{L}(\zeta).$ (3.17)
Proof.
According to Lemma 3.3, (3.16) and the Sobolev embedding theorem, for any $1\geq\varepsilon>0$ and $\zeta\in E$, there exists a real number $t^{*}>0$ such that
$\|z_{0}^{\zeta}(t^{*})-(m_{L}-\psi)\|_{E}+\|z_{0}^{\zeta}(t^{*})-(m_{L}-\psi)\|_{H}\leq\varepsilon,$
and
$\|\Delta z_{0}^{\zeta}(t^{*},\cdot)+F(z_{0}^{\zeta}(t^{*}))(\cdot)\|_{H}\leq\varepsilon.$
Define
$v(s)=(1-s)(m_{L}-\psi)+s\cdot z_{0}^{\zeta}(t^{*})=(m_{L}-\psi)+s(z_{0}^{\zeta}(t^{*})-(m_{L}-\psi)),\;s\in[0,1].$
By definition, $v$ is a path in $C([0,1];E)$ that satisfies $v(0)=m_{L}-\psi$ and $v(1)=z_{0}^{\zeta}(t^{*})$.
Note that, under our assumptions, the following more explicit expression for $I_{T}^{x}$ can be read as:
$I_{T}^{x}(z)=\frac{1}{2}\int_{0}^{T}\left\|\frac{\partial_{t}z-\Delta z(t)-F(z(t))}{G(t,z(t))}\right\|_{H}^{2}\;dt.$
Thus,
$g_{0}\cdot I_{1}^{m_{L}-\psi}(v)$ $\leq\frac{1}{2}\int_{0}^{1}\|\partial_{s}v-\Delta v(s)-F(v(s))\|_{H}^{2}\;ds$
$\leq\frac{1}{2}\|z_{0}^{\zeta}(t^{*})-(m_{L}-\psi)\|_{H}^{2}+\frac{1}{2}\int_{0}^{1}\|\Delta v(s)+F(v(s))\|_{H}^{2}\;ds$
$\leq\frac{\varepsilon}{2}+\int_{0}^{1}\|\Delta(m_{L}-\psi)+F(m_{L}-\psi)\|_{H}^{2}+\|s[\Delta z_{0}^{\zeta}(t^{*})-\Delta(m_{L}-\psi)]\|_{H}^{2}$
$\ \ \ \ +\|F(v(s))-F(m_{L}-\psi)\|_{H}^{2}\;ds$
$=\frac{\varepsilon}{2}+\int_{0}^{1}\left[\|s(\Delta z_{0}^{\zeta}(t^{*})-\Delta(m_{L}-\psi))\|_{H}^{2}+\|F(v(s))-F(m_{L}-\psi)\|_{H}^{2}\right]\;ds.$
Since $F(\cdot)$ is locally Lipschitz continuous,
$\|F(z_{0}^{\zeta}(t^{*}))-F(m_{L}-\psi)\|_{H}\leq CL\varepsilon,$
and
$\int_{0}^{1}\|F(v(s))-F(m_{L}-\psi)\|_{H}^{2}\;ds\leq C^{2}L^{2}\varepsilon^{2}.$
Here, the last inequality is due to $\|v(s)-(m_{L}-\psi)\|_{E}\leq\varepsilon$, $s\in[0,1]$. Then,
$\|\Delta z_{0}^{\zeta}(t^{*})-\Delta(m_{L}-\psi)\|_{H}$ $\leq\|\Delta z_{0}^{\zeta}(t^{*})+F(z_{0}^{\zeta}(t^{*}))-[\Delta(m_{L}-\psi)+F(m_{L}-\psi)]\|_{H}$
$\ \ \ \ +\|F(z_{0}^{\zeta}(t^{*}))-F(m_{L}-\psi)\|_{H}$
$\leq\|\Delta z_{0}^{\zeta}+F(z_{0}^{\zeta}(t^{*}))\|_{H}+CL\varepsilon\leq(CL+1)\varepsilon.$
According to the above discussions,
$g_{0}\cdot I_{1}^{m_{L}-\psi}\leq(CL+1)^{2}\varepsilon.$ (3.18)
Define
$\tilde{v}(s)=z_{0}^{\zeta}(t^{*}+1-s),\;s\in[1,t^{*}+1].$
We have $\tilde{v}(1)=z_{0}^{\zeta}(t^{*})$ and $\tilde{v}(t^{*}+1)=\zeta$. Moreover,
$\partial_{s}\tilde{v}=-\Delta z_{0}^{\zeta}(t^{*}+1-s)-F(z_{0}^{\zeta}(t^{*}+1-s)).$
Thus,
$g_{0}\cdot I_{[1,t^{*}+1]}^{z_{0}^{\zeta}(t^{*})}(\tilde{v})\leq 2\int_{0}^{t^{*}}\|\Delta z_{0}^{\zeta}(s)+F(z_{0}^{\zeta}(s))\|_{H}^{2}\;ds\leq 2\mathbf{E}_{L}^{*}(\zeta),$ (3.19)
where the last inequality is by (3.16). Therefore, for any $\zeta\in H^{1}(-L,L)$, there exists $z\in C([0,t^{*}+1];E)$ defined as
$z(t)=\begin{cases}v(t),\;t\in[0,1],\\ \tilde{v}(t),\;t\in[1,t^{*}+1],\end{cases}$
such that $z$ satisfies $z(0)=m_{L}-\psi$ and $z(t^{*}+1)=\zeta$. Moreover, by (3.18) and (3.19), for any $\zeta\in H^{1}(-L,L)$, $\varepsilon>0$, there exists $t^{*}>0$ such that
$g_{0}\cdot I_{t^{*}+1}^{m_{l}-\psi}(z)\leq g_{0}\cdot I_{1}^{m_{L}-\psi}(v)+g_{0}\cdot I_{[1,t^{*}+1]}^{z_{0}^{\zeta}(t^{*})}(\tilde{v})\leq 2\mathbf{E}_{L}^{*}(\zeta)+C(L+1)\varepsilon.$
This is then followed by
$g_{0}\cdot U_{L}\leq 2\mathbf{E}^{*}_{L}(\zeta).$
3.2 Compactness of level sets of the quasi-potential
In this section, we will show the compactness of level sets of quasi-potential. For any $t_{1}<t_{2}$ and $z\in C([t_{1},t_{2}];E)$, define
$I_{t_{1},t_{2}}(z):=\frac{1}{2}\inf\{\|f\|_{L^{2}(t_{1},t_{2};H)}^{2};f\in L^{2}(t_{1},t_{2};H),z=z_{f}\},$
where $z_{f}$ is the solution of the skeleton equation (2.3) in $t_{1},t_{2}]$ with $z_{f}(t_{1})=z(t_{1})$. We write $I_{-t}$ when $t_{2}=0$ and $t_{1}=-t<0$.
According to the proof of Theorem 5.1 in [cerrai2004large], for any compact set $\Lambda\in E$ the level set
$K_{\Lambda,t_{1},t_{2}}(r):=\{z\in C([t_{1},t_{2}];E),z(t_{1})\in\Lambda,I_{t_{1},t_{2}}(z)\leq r\}$
is compact.
Similarly, for $z\in C((-\infty,0];E)$, define
$I_{-\infty}(z):=\frac{1}{2}\inf\{\|f\|_{L^{2}(-\infty,0;H)}^{2};f\in L^{2}(-\infty,0;H),z=z_{f}\}=\sup_{t\geq 0}I_{-t}(z),$
and for $r\geq 0$
$K_{-\infty}(r):=\{z\in C((-\infty,0];E),I_{-\infty}(z)\leq r,\lim_{t\rightarrow+\infty}\|z(-t)-(m_{L}-\psi)\|_{E}=0\}.$
The following lemma is needed later in this paper.
Lemma 3.5.
For any $r\geq 0$, if
$\liminf_{t\rightarrow+\infty}\|z(-t)\|_{E}<\infty,I_{-\infty}(z)\leq r$
then
$\sup_{t\geq 0}\mathbf{E}^{*}_{L}(z(-t))\leq C_{r},$
where $C_{r}>0$ is a constant depending on $r$.
Proof.
According to the assumption of the lemma, there exist a sequence $t_{n}\nearrow+\infty$ and a constant $C>0$ such that $\|z(-t_{n})\|_{E}\leq C$, $n\in\mathbb{N}$ and
$f\in L^{2}(-\infty,0;H),\;\sup_{n\in\mathbb{N}}I_{-t_{n}}(z)=\frac{1}{2}\|f\|_{L^{2}(-\infty,0;H)}^{2}\leq r,$
where $z$ is the solution of equation (2.3) with $z(-t_{n},\cdot)=z(-t_{n})$. Due to (3.12),
$\sup_{t\in[-t_{n},0],n\in\mathbb{N}}\|z(t)\|_{E}<\infty.$ (3.20)
Then, by standard estimates,
$\frac{1}{2}\frac{d}{dt}\|z(t)\|_{H}^{2}$ $=\langle\Delta z(t),z(t)\rangle_{H}+\langle F(z)(t),z(t)\rangle+\langle G(t,z(t))f(t),z(t)\rangle$
$\leq-\|z(t)\|_{H^{1}}^{2}-\kappa\|z(t)\|_{H}^{2}+C\|f(t)\|_{H}^{2}+C_{\kappa}.$
Note that $\|f\|_{L^{2}([-t_{n},0];H)}^{2}=2I_{t_{n}}(z_{n})\leq 2r$, by Grönwall’s inequality, we have for $t\in[-t_{n},0]$
$\|z(t)\|_{H}^{2}+2\int_{-t_{n}}^{t}\|z(s)\|_{H^{1}}^{2}\;ds\leq C(t+t_{n})(\|z(-t_{n})\|_{H}^{2}+2r+1).$
Thus, for any $n\in\mathbb{N}$ there exists $t^{\prime}\in[-t_{n},-t_{n}+1]$ such that $\|z(t^{\prime})\|_{H^{1}}^{2}\leq C(\|z(-t_{n})\|_{H}^{2}+2r+1)$. From (1.7) and the Sobolev embedding, $\mathbf{E}^{*}_{L}(z(t^{\prime}))\leq C_{r}$. Then, by (3.11), for any $t\geq t^{\prime}$
$\mathbf{E}^{*}_{L}(z(t))\leq\mathbf{E}^{*}_{L}(z(t^{\prime}))+2C\cdot r\leq C_{r}.$
The proof is finished by taking $t_{n}$ to -$\infty$. ∎
Proposition 3.6.
For any $r\geq 0$, the set $K_{-\infty}(r)$ is compact in $C((-\infty,0];E)$.
Before we prove this proposition, we need the following lemma.
Lemma 3.7.
Suppose $\|z(0)-(m_{L}-\psi)\|_{E}\geq\beta$ and $\mathbf{E}_{L}^{*}(z(0))\leq C_{r}$ for some $\beta,C_{r}>0$, then there exists $t_{0}>0$, such that
$I_{t_{0}}(z)\geq\beta.$
Proof.
We will show that there exists $t_{0}$ such that
$r_{t_{0}}:=\inf\{I_{t_{0}}(z):z\in C([0,t_{0}];E),\mathbf{E}^{*}_{L}(z(0))\leq C_{r},\|z(t_{0})-(m_{L}-\psi)\|_{E}\geq\beta\}>0.$
In fact, if $r_{t_{0}}=0$, then there exists a sequence
$\{z_{i}\in C([0,t_{0}];E),\mathbf{E}^{*}_{L}(z_{i}(0))\leq C_{r},\|z_{i}(t_{0})-(m_{L}-\psi)\|_{E}\geq\beta\},i\in\mathbb{N},$ (3.21)
such that
$\lim_{i\rightarrow\infty}I_{t_{0}}(z_{i}(t))=0.$ (3.22)
This implies that there exists $f_{i}$ with $\lim_{i\rightarrow\infty}\|f_{i}\|_{L^{2}(0,t_{0};H)}=0$ such that $z_{i}$ is the solution of the skeleton equation
$\partial_{t}z_{i}(t)=\Delta z_{i}(t)+F(z_{i}(t))+G(t,z_{i}(t))f_{i}(t).$
According to (3.12),
$\sup_{t\geq 0,i\in\mathbb{N}}\|z_{i}\|_{E}<\infty.$ (3.23)
Since the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$ ([bertini2008dobrushin]), there exist $x\in E$ and a subsequence still denoted by $z_{i}(0)$ such that
$\lim_{i\rightarrow\infty}\|z_{i}(0)-x\|_{H}\leq L\cdot\lim_{i\rightarrow\infty}\|z_{i}(0)-x\|_{E}=0.$
Then, by standard estimates,
$\frac{1}{2}\frac{d}{dt}\|z_{i}(t)-z_{0}^{x}(t)\|_{H}^{2}$ $\leq\langle\Delta(z_{i}(t)-z_{0}^{x}(t)),z_{i}(t)-z_{0}^{x}(t)\rangle_{H}+\langle F(z_{i}(t))-F(z_{0}^{x}(t)),z_{i}(t)-z_{0}^{x}(t)\rangle$
$+\langle G(t,z_{i}(t))f_{i}(t),z_{i}(t)-z_{0}^{x}(t)\rangle$
$\leq-\|z_{i}(t)-z_{0}^{x}(t)\|_{H^{1}}^{2}+C\|z_{i}(t)-z_{0}^{x}(t)\|_{H}^{2}+C\|f_{i}(t)\|_{H}^{2}.$
By Grönwall’s inequality
$\|z_{i}(t)-z_{0}^{x}(t)\|_{H}^{2}+2\int_{0}^{t}\|z_{i}(s)-z_{0}^{x}(s)\|_{H^{1}}^{2}\;ds\leq C_{t}(\|z_{i}(0)-x\|_{H}^{2}+\|f_{i}\|_{L^{2}([0,t];H)}^{2}).$
Then, there exists $t\in[\frac{1}{2}t_{0},t_{0}]$ such that
$\|z_{i}(t)-z_{0}^{x}(t)\|_{H^{1}}^{2}\leq\frac{C_{t_{0}}}{t_{0}}(\|z_{i}(0)-x\|_{H}^{2}+\|f_{i}\|_{L^{2}(0,t_{0};H)}^{2}).$ (3.24)
Thanks to Lemma 3.3, for any $\eta>0$, one can choose $t_{0}$ sufficiently large so that $\mathbf{E}^{*}_{L}(z_{0}^{x}(t))\leq\frac{\eta}{4}$. Thus, by the Sobolev inequality and (3.24), there exists $N$ such that for any $N\leq i\in\mathbb{N}$,
$|\mathbf{E}^{*}_{L}(z_{i}(t))-\mathbf{E}^{*}_{L}(z_{0}^{x}(t))|\leq C\|z_{i}(t)-z_{0}^{x}(t)\|_{H^{1}}^{2}\leq\frac{\eta}{4}.$
Then,
$\mathbf{E}^{*}_{L}(z_{i}(t))\leq\mathbf{E}^{*}_{L}(z_{0}^{x}(t))+\frac{\eta}{4}\leq\frac{\eta}{2}.$
Furthermore, by (3.11) and (3.23),
$\mathbf{E}^{*}_{L}(z_{i}(t_{0}))$ $\leq\mathbf{E}^{*}_{L}(z_{i}(t))+C(\sup_{s\geq 0,i\in\mathbb{N}}\|z_{i}(s)\|_{E}^{2}+1)\cdot I_{t_{0}}(z_{i})$
$\leq\mathbf{E}^{*}_{L}(z_{i}(t))+CI_{t_{0}}(z_{i}).$
By (3.22), for any $\eta>0$, there exists $N$ such that for all $i\geq N$, $CI_{t_{0}}(z_{i})\leq\frac{\eta}{2}$. Therefore, one obtains that for $i$ sufficiently large,
$\mathbf{E}^{*}_{L}(z_{i}(t_{0}))\leq\eta.$
Moreover, by Proposition 3.1, for any $\beta>0$, one can choose $\eta$ sufficiently small such that
$\|z_{i}(t_{0})-(m_{L}-\psi)\|_{E}\leq\frac{\beta}{2}$
which contradicts the assumption that $\|z_{i}(t_{0})-(m_{L}-\psi)\|_{E}\geq\beta$ in (3.21). ∎
Proof of Proposition 3.6.
Fixing $r\geq 0$, we need to show that for any sequence $\{z_{n}\}\subset K_{-\infty}(r)$, there exists a subsequence $\{z_{n_{k}}\}$ converging in $C((-\infty,0];E)$ to some $\hat{z}\in K_{-\infty}(r)$. According to Lemma 3.5, for any $k\in\mathbb{N}$, the restriction of $z_{n}$ to the interval [k,0][-k,0] belongs to $K_{\Lambda,-k,0}$, where
$\Lambda:=\{x\in E;\mathbf{E}^{*}_{L}(x)\leq C_{r}\},$
and $C_{r}$ is given in Lemma 3.5. As the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$, we have $K_{\Lambda,-k,0}$ is compact in C([k,0];E)C([-k,0];E) for any $k\in\mathbb{N}$.
We choose a subsequence following Helly’s diagonal procedure. Taking $k=1$ we can find $\{z_{n_{1}}\}\subset\{z_{n}\}$ and $\hat{z}_{1}\in C([-1,0];E)$ such that $z_{n_{1}|_{[-1,0]}}$ converges to $\hat{z}_{1}$ in C([1,0];E)C([-1,0];E). With the same arguments, we can find a subsequence $\{z_{n_{2}}\}\subset\{z_{n_{1}}\}$ and $\hat{z}_{2}\in C([-2,0];E)$ such that $z_{n_{2}|_{[-2,0]}}$ converges to $\hat{z}_{2}$ in C([2,0];E)C([-2,0];E). It is obvious that $\hat{z}_{2}|_{[-1,0]}=\hat{z}_{1}$. By this procedure, we can find a subsequence $\{z_{n^{\prime}}\}\subseteq\{z_{n}\}$ converging to some $\hat{z}$ in $C((-\infty,0];E)$. Thanks to Theorem 5.1 of [cerrai2004large], $I_{-k}(\cdot)$ is lower semi-continuous and we have $I_{-k}(\hat{z})\leq r$ for any $k\geq 0$ which implies $I_{-\infty}(\hat{z})\leq r$. Moreover, by Lemma 3.5, we have that
$\sup_{t\geq 0}\mathbf{E}^{*}_{L}(\hat{z}(-t))\leq C_{r}.$
To show that $\hat{z}\in K_{-\infty}(r)$, it remains to prove that
$\lim_{t\rightarrow-\infty}\|\hat{z}(t)-(m_{L}-\psi)\|_{E}=0.$
If this does not hold, there exist a constant $\beta>0$ and a sequence $t_{n}\nearrow+\infty$ such that $\|\hat{z}(-t_{n})-(m_{L}-\psi)\|_{E}\geq\beta$, for any $n\in\mathbb{N}$. Then, by Lemma 3.7, there exists $t_{0}>0$ such that $I_{-t_{n}-t_{0},-t_{n}}(\hat{z})\geq r_{t_{0}}>0$, for $n\in\mathbb{N}$. Therefore, if we choose a subsequence $\{t_{n_{k}}\}\subset\{t_{n}\}$ such that $t_{n_{k+1}}\geq t_{n_{k}}+t_{0}$. Then, for any $m\in\mathbb{N}$, we have
$I_{-\infty}(\hat{z})\geq\sum_{k=1}^{m}I_{-t_{n_{k+1}},-t_{n_{k}}}(\hat{z})\geq\sum_{k=1}^{m}I_{-t_{n_{k}}-t_{0},-t_{n_{k}}}(\hat{z})\geq mr_{t_{0}}.$
[…]
We have the following characterisation of $U_{L}$. For simplicity of presentation, we denote
$I_{-\infty}^{\inf}(x)=\inf\{I_{-\infty}(z);z\in C((-\infty,0];E),\lim_{t\rightarrow+\infty}\|z(-t)-(m_{L}-\psi)\|_{E}=0,z(0)=x\}.$
If the infimum is attainable, we write $I_{-\infty}^{\inf}(x)$ as $I_{-\infty}^{\min}(x)$.
Proposition 3.8.
Under Assumption 1, for any $x\in E$
$U_{L}(x)=I_{-\infty}^{\min}(x).$
Proof.
Let $T>0$ be fixed and $z\in C([-T,0];E)$ such that $z(-T)=m_{L}-\psi$, $z(0)=x$ and $I_{-T}(z)<\infty$. Set
$\bar{z}(t):=\begin{cases}z(t),&t\in[-T,0],\\ m_{L}-\psi,&t\in(-\infty,-T].\end{cases}$
Obviously, $\bar{z}\in C((-\infty,0];E)$, $\bar{z}(0)=x$ and $\lim_{t\rightarrow+\infty}\|\bar{z}(-t)-(m_{L}-\psi)\|_{E}=0$. Since $m_{L}-\psi$ is the solution for equation (3.2), we have
$I_{-\infty}(\bar{z})=I_{-T}(\bar{z})+I_{-\infty,-T}(\bar{z})=I_{-T}(z).$
In particular, this implies that
$I_{-\infty}^{\inf}(x)\leq I_{-T}(z).$
Taking the infimum over all $z\in C([-T,0];E)$ and $T>0$, we get
$I_{-\infty}^{\inf}(x)\leq U_{L}(x).$
Next, we will prove the opposite. If
$I_{-\infty}^{\inf}(x)=+\infty,$
there is nothing to prove. Therefore, we only need to consider $z\in K_{-\infty}(r)$, $z(0)=x$, for some $r>0$. For such $z$, there exists a sequence $t_{n}\nearrow+\infty,n\in\mathbb{N}$ such that
$\lim_{n\rightarrow+\infty}\|z(-t_{n})-(m_{L}-\psi)\|_{H}\leq L\cdot\left(\lim_{n\rightarrow+\infty}\|z(-t_{n})-(m_{L}-\psi)\|_{E}\right)=0.$
and
$f\in L^{2}(-\infty,0;H),\ \sup_{n\in\mathbb{N}}I_{-t_{n}}=\frac{1}{2}\|f\|_{L^{2}(-\infty,0;H)}^{2}\leq r,$
where $z$ is the solution of equation (2.3), starting from time $t_{n}$. Since $f\in L^{2}(-\infty,0;H)$, for any $\varepsilon>0$, there exists $T_{\varepsilon}\leq 0$ such that $\frac{1}{2}\|f\|_{L^{2}(-\infty,T_{\varepsilon};H)}^{2}\leq\varepsilon$. Denote $z_{f}^{-t_{n},m_{L}-\psi}$ as the solution of equation (2.3) with $z_{f}^{-t_{n},m_{L}-\psi}(-t_{n})=m_{L}-\psi$. Since $m_{L}-\psi$ is the equilibrium of (2.3) with $f=0$, we have $z_{0}^{-t_{n},m_{L}-\psi}(t)=m_{L}-\psi$ for every $t\in[-t_{n},0]$. Then, by (3.24), for $n\in\mathbb{N}$ sufficiently large there exists $T_{\varepsilon}\in[-t_{n},-t_{n}+1]$ such that
$\|z(T_{\varepsilon})-(m_{L}-\psi)\|_{H^{1}}^{2}\leq C(\|z(-t_{n})-(m_{L}-\psi)\|_{H}^{2}+2\varepsilon)\leq C\varepsilon.$
According to discussions in Proposition 3.4, there exists a path $v_{\varepsilon}$ such that $v_{\varepsilon}(T_{\varepsilon}-1)=m_{L}-\psi$, $v_{\varepsilon}(T_{\varepsilon})=z(T_{\varepsilon})$ and $I_{T_{\varepsilon}-1,T_{\varepsilon}}(v_{\varepsilon})\leq C\epsilon$. We set
$\bar{v}_{\varepsilon}(t):=\begin{cases}v_{\varepsilon}(t),&t\in[T_{\varepsilon}-1,T_{\varepsilon}],\\ z(t),&t\in[T_{\varepsilon},0].\end{cases}$
Then, $\bar{v}_{\varepsilon}(T_{\varepsilon}-1)=m_{L}-\psi$, $\bar{v}_{\varepsilon}(0)=x$, and
$I_{T_{\varepsilon}-1}(\bar{v}_{\varepsilon})\leq I_{-\infty}(z)+C\varepsilon.$
By the definition of $U_{L}$,
$U_{L}(x)\leq I_{T_{\varepsilon}-1}(\bar{v}_{\varepsilon}).$
This then implies
$U_{L}(x)\leq I_{-\infty}(z)+\varepsilon.$
Therefore, by taking the infimum over $\{z\in C((-\infty,0];E),\lim_{t\rightarrow+\infty}\|z(-t)-(m_{L}-\psi)\|_{E}=0,z(0)=x\}$, and since $\epsilon>0$ can be arbitrarily small, we get
$U_{L}(x)\leq I_{-\infty}^{\inf}(x).$
Finally, by Proposition 3.6, $K_{-\infty}(r)$ is compact. Hence, $I_{-\infty}^{\inf}(x)$ is indeed $I_{-\infty}^{\min}(x)$. The proof is complete. ∎
The characterisation of $U_{L}$ allows us to enhance the compactness of the level sets of $I_{-\infty}(\cdot)$ on $C((-\infty,0];E)$ to the compactness of the level sets of $U_{L}(\cdot)$ on $E$. This is the main result of this section and a key result in the theory of large deviations for invariant measures.
Theorem 3.9.
Under Assumption 1, for any $r\geq 0$ the level set
$\Phi(r):=\{x\in E:U_{L}(x)\leq r\}$
is compact in $E$.
Proof.
Let $\{x_{n}\}\subset\Phi_{L}(r)$. According to Proposition 3.8, for each $n\in\mathbb{N}$ we can find $z_{n}\in C((-\infty,0];E)$ with $z_{n}(0)=x_{n}$ and
$\lim_{t\rightarrow+\infty}\|z_{n}(-t)-(m_{L}-\psi)\|_{E}=0,$
such that $U_{L}(x_{n})=I_{-\infty}(z_{n})$. Since $U_{L}(x_{n})\leq r$, we have $\{z_{n}\}\subset K_{-\infty}(r)$. By Proposition 3.6, $K_{-\infty}(r)$ is compact. So, there exists a subsequence $\{z_{n_{k}}\}\subset\{z_{n}\}$ such that $z_{n_{k}}\rightarrow z\in K_{-\infty}(r)$ in $C((-\infty,0];E)$. In particular, $x_{n_{k}}=z_{n_{k}}(0)\rightarrow z(0)$ in $E$. Thanks to Proposition 3.8, we have $U_{L}(z(0))\leq I_{-\infty}(z)\leq r$. That states $z(0)\in\Phi_{L}(r)$. The proof is complete. ∎
4 LDP for invariant measures
In this section, as the main result of this paper, we will prove the LDP for $\{\mu_{\epsilon}\}_{\epsilon>0}$. This is indicated by the LDP of $\{\bar{\mu}_{\epsilon}\}_{\epsilon>0}$, whose relation with $\{\mu_{\epsilon}\}_{\epsilon>0}$ is given by (1.5).
Theorem 4.1.
Under Assumption 1, the family of invariant measures $\{\bar{\mu}_{\epsilon}\}_{\epsilon}$ satisfies the large deviation principle in $E$ with a good rate function $U_{L}$.
4.1 Exponential estimates
The essential part of the proof is an exponential estimate for $\bar{\mu}_{\epsilon}$ over a compact subset of $E$. We fix $0<k^{\star}<\frac{1}{2}$ and an even number $p^{\star}$ such that $k^{\star}p^{\star}>1$ and $p^{\star}>6$. Then, we can find some $0<\alpha<1/4$ that depend on $k^{\star},p^{\star}$ such that
$(\alpha-1-\frac{k^{\star}}{2})\frac{p^{\star}}{p^{\star}-1}>-1.$ (4.1)
By the Sobolev embedding theorem, $W^{k^{\star},p^{\star}}(-L,L)$ is compactly embedded into $E$. We will first prove the following proposition.
Proposition 4.2.
For any $r>0$, there exist $\epsilon_{r}>0$ and $R_{r}>0$ such that
$\bar{\mu}_{\epsilon}(B_{W^{k^{\star},p^{\star}}(-L,L)}^{c}(0,R_{r}))\leq\exp(-\frac{r}{\epsilon}),\;\epsilon<\epsilon_{r}.$ (4.2)
Proof.
Rewrite equation (1.3) in a different form for $\bar{u}_{\epsilon}(0)=0$ and $\lambda>0$
$\partial_{t}\bar{u}_{\epsilon}^{0}(t,\xi)=(\Delta-\lambda)\bar{u}_{\epsilon}(t,\xi)+F(\bar{u}_{\epsilon}^{0}(t))(\xi)+\lambda\bar{u}_{\epsilon}^{0}(t,\xi)+\epsilon^{\frac{1}{2}}G(t,\bar{u}_{\epsilon}^{0}(t))\partial_{t}W(t,\xi).$ (4.3)
The mild solution $\bar{u}_{\epsilon}^{0}(t)$ is represented as
$\bar{u}_{\epsilon}^{0}(t)=\int_{0}^{t}S(t-s)e^{-\lambda(t-s)}[F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)]\,ds+\epsilon^{\frac{1}{2}}\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t),$
where,
$\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t):=\int_{0}^{t}S(t-s)e^{-\lambda(t-s)}G(s,\bar{u}_{\epsilon}^{0}(s))\,dW(s).$
Then,
$\begin{split}&\ \ \ \ \|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}=\|(-\Delta)^{\frac{k^{\star}}{2}}\bar{u}_{\epsilon}^{0}(t)\|_{L^{p^{\star}}}\\ &\leq\int_{0}^{t}\left\|(-\Delta)^{\frac{k^{\star}}{2}}S(t-s)e^{-\lambda(t-s)}[F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)]\right\|_{L^{p^{\star}}}\,ds\\ &\ \ \ \ +\epsilon^{\frac{1}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\\ &\leq\int_{0}^{t}(t-s)^{-\frac{k^{\star}}{2}}e^{-\lambda(t-s)}\;ds\cdot\sup_{s\in[0,t]}\|F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)\|_{L^{p^{\star}}}\\ &\ \ \ \ +\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\\ &\leq C\sup_{s\in[0,t]}\|F(\bar{u}_{\epsilon}^{0}(s))+\lambda\bar{u}_{\epsilon}^{0}(s)\|_{L^{p^{\star}}}+\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\\ &\leq C\sup_{s\in[0,t]}\left(\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3}+\epsilon^{\frac{3}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3}+\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}+1\right).\end{split}$ (4.4)
Here,
$Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t):=\bar{u}_{\epsilon}^{0}(t)-\epsilon^{\frac{1}{2}}\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)$ (4.5)
which satisfies the equation
$\begin{cases}\partial_{t}Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t)=(\Delta-\lambda)Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t)+F(\bar{u}_{\epsilon}^{0}(t))+\lambda\bar{u}_{\epsilon}^{0}(t),\\ Y_{\lambda}(\bar{u}_{\epsilon}^{0})(0)=0.\end{cases}$ (4.6)
Next, we need to estimate $\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})\|_{L^{3p^{\star}}}^{3p^{\star}}$. Multiplying both sides of equation (4.6) by $3p^{\star}Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-1}(t)$ and integrating with respect to $\xi$ in [L,L][-L,L], since $p^{\star}>0$ is chosen to be an even number, by Young’s inequality, there exists $0<\sigma<1$ such that
$\frac{d\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})\|_{L^{3p^{\star}}}^{3p^{\star}}}{dt}$ $=3p^{\star}\int_{-L}^{L}(\Delta-\lambda)Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t,\xi)\cdot Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-1}(t,\xi)\;d\xi$
$\ \ \ \ +3p^{\star}\int_{-L}^{L}[F(\bar{u}_{\epsilon}^{0}(t,\xi))+\lambda\bar{u}_{\epsilon}^{0}(t,\xi)]\cdot Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-1}(t,\xi)\;d\xi$
$=-3p^{\star}(3p^{\star}-1)\int_{-L}^{L}[\nabla Y_{\lambda}(\bar{u}_{\epsilon}^{0})(t,\xi)]^{2}Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-2}(t,\xi)\;d\xi$
$\ \ \ \ -\int_{-L}^{L}\lambda Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}}(t,\xi)\;d\xi$
$\ \ \ \ +3p^{\star}\int_{-L}^{L}[F(\bar{u}_{\epsilon}^{0}(t,\xi))+\lambda\bar{u}_{\epsilon}^{0}(t,\xi)]\cdot Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}-1}(t,\xi)\;d\xi$
$\leq 3p^{\star}\int_{-L}^{L}\left[-(1-\sigma)Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}+2}(t,\xi)+C_{\sigma}\epsilon^{\frac{3p^{\star}+2}{2}}\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}+2}(t,\xi)+C_{1}\right]\;d\xi$
$\leq\int_{-L}^{L}\left[-C_{2}Y_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}}(t,\xi)+C_{3}[\epsilon^{\frac{1}{2}}\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})^{3p^{\star}+2}(t,\xi)+1]\right]\;d\xi.$
By Grönwall’s inequality, this implies that
$\sup_{s\in[0,t]}\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3p^{\star}}\leq C\left(\sup_{s\in[0,t]}\epsilon^{\frac{3p^{\star}+2}{2}}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}+2}}^{3p^{\star}+2}+1\right).$ (4.7)
Thus, by (4.4), the Sobolev inequality and (4.7),
$\mathbb{P}(\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r})$ $\leq\mathbb{P}\left(\sup_{s\in[0,t]}\|Y_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}^{3p^{\star}}\geq\left(\frac{R_{r}/3-C}{C}\right)^{p^{\star}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{3p^{\star}}}\geq\left(\frac{R_{r}}{3C}\right)^{\frac{1}{3}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\frac{R_{r}}{3}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\left[\frac{\left(\frac{R_{r}/3-C}{C}\right)^{p^{\star}}-C}{C}\right]^{\frac{1}{3p^{\star}+2}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\left(\frac{R_{r}}{3C}\right)^{\frac{1}{3}}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\frac{R_{r}}{3}\right)$
$\leq 3\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right).$
Here, $\tilde{R}_{r}$ is a constant derived from $R_{r}$. Hence,
$\bar{\mu}_{\epsilon}(B_{W^{k^{\star},p^{\star}}(-L,L)}^{c}(0,R_{r}))$ $=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\mathbb{P}(\|\bar{u}_{\epsilon}^{0}(t)\|_{W^{k^{\star},p^{\star}}}\geq R_{r})\;dt$
$\leq\lim_{T\rightarrow\infty}\frac{3}{T}\int_{0}^{T}\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)\;dt.$
To finish the proof, it is sufficient to show the following result.
Lemma 4.3.
For any $r>0$, there exist $\epsilon_{r}>0$ and $\tilde{R}_{r}>0$ such that
$\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)\;dt\leq\exp(-\frac{r}{\epsilon}),\;\epsilon<\epsilon_{r}.$ (4.8)
Proof.
Notice that
$\ \ \ \ \mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)$
$=\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)$
$\ \ \ \ +\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}>\bar{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq\bar{R}_{r}\right)$
$\ \ \ \ +\mathbb{P}\left(\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\geq\bar{R}_{r}\right).$
By the factorisation formula,
$\gamma_{\lambda}(v)(t)=C_{\alpha}\int_{0}^{t}(t-s)^{\alpha-1}S(t-s)e^{-\lambda(t-s)}\Gamma^{\alpha}(v)(s)\;ds,$
where
$\Gamma^{\alpha}(v)(s)=\int_{0}^{s}(s-r)^{-\alpha}S(s-r)e^{-\lambda(s-r)}G(v(r))\;dW(r).$
Then, by (4.1)
$\begin{split}\|\gamma_{\lambda}(v)(t)\|_{W^{k^{\star},p^{\star}}}^{p^{\star}}&\leq C_{\alpha}^{p^{\star}}\left(\int_{0}^{t}(t-s)^{\alpha-1-\frac{k^{\star}}{2}}e^{-\lambda(t-s)}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}\;ds\right)^{p^{\star}}\\ &\leq C_{\alpha}^{p^{\star}}\left(\int_{0}^{t}(t-s)^{(\alpha-1-\frac{k^{\star}}{2})\frac{p^{\star}}{p^{\star}-1}}e^{-\frac{\lambda}{2}(t-s)\frac{p^{\star}}{p^{\star}-1}}\;ds\right)^{(p^{\star}-1)}\\ &\ \ \ \ \times\left(\int_{0}^{t}e^{-\frac{\lambda}{2}(t-s)p^{\star}}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}\;ds\right)\\ &\leq C\times\left(\int_{0}^{t}e^{-\frac{\lambda}{2}(t-s)p^{\star}}\;ds\right)\times\sup_{s\in[0,t]}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}\\ &\leq C\sup_{s\in[0,t]}\|\Gamma^{\alpha}(v)(s)\|_{L^{p^{\star}}}^{p^{\star}}.\end{split}$ (4.9)
Thus, for any $\|v\|_{C([0,t];E)}\leq\bar{R}_{r}$,
$\sup_{s\in[0,t]}\|\gamma_{\lambda}(v)(s)\|_{W^{k^{\star},p^{\star}}}\leq C\sup_{s\in[0,t]}\left\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\right\|_{L^{p^{\star}}},$ (4.10)
where
$\Gamma^{\alpha}_{\bar{R}_{r}}(t)=\int_{0}^{t}(t-s)^{-\alpha}S(t-s)e^{-\lambda(t-s)}(1+\bar{R}_{r})\;dW(s).$
Denote
$e_{k}(\xi)=\begin{cases}\frac{\sin(\frac{k\pi\xi}{2L})}{\sqrt{L}},\;k\in\mathbb{N}\;\text{and is even}\\ \frac{\cos(\frac{k\pi\xi}{2L})}{\sqrt{L}},\;k\in\mathbb{N}\;\text{and is odd.}\end{cases}$ (4.11)
Clearly, $e_{k}$ is the orthogonal basis of $H$. Thus, for any $t\geq 0$, $\xi\in[-L,L]$
$\mathbb{E}|\Gamma^{\alpha}_{\bar{R}_{r}}(t,\xi)|^{2}$ $\leq\sum_{k\in\mathbb{N}}\int_{0}^{t}s^{-2\alpha}e^{-2(\frac{L^{-1}k\pi}{2})^{2}s}(1+\bar{R}_{r})^{2}\;ds\cdot e_{k}^{2}(\xi)$
$\leq C\sum_{k\in\mathbb{N}}\frac{1}{(L^{-1}k\pi)^{2(1-2\alpha)}}\cdot(1+\bar{R}_{r})^{2}$
$\leq C_{\alpha}(1+\bar{R}_{r})^{2},$
where $C_{\alpha}>0$ depends on $\alpha$ and does not depend on $t$. Therefore, by the Burkholder–Davis–Gundy inequality, for any $t>0$
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\Gamma^{\alpha}_{\bar{R}_{r}}(s)\|_{L^{p^{\star}}}^{p^{\star}}\right]$ $\leq\mathbb{E}\int_{-L}^{L}\sup_{s\leq t}|\Gamma^{\alpha}_{\bar{R}_{r}}(s,\xi)|^{p^{\star}}\;d\xi$
$\leq\int_{-L}^{L}C_{p^{\star}}(\mathbb{E}|\Gamma^{\alpha}_{\bar{R}_{r}}(t,\xi)|^{2})^{\frac{p^{\star}}{2}}\;d\xi$
$\leq C_{p^{\star}}C_{\alpha}^{\frac{p^{\star}}{2}}(1+\bar{R}_{r})^{p^{\star}}.$
Taking $t\rightarrow\infty$ on the left hand side of the above inequality and by the monotone convergence theorem, we obtain that
$\mathbb{E}\left[\sup_{t\geq 0}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}^{p^{\star}}\right]\leq C_{\alpha,p^{\star}}(1+\bar{R}_{r})^{p^{\star}}.$
By Chebyshev’s inequality, for $\delta>0$
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}\geq\delta\right)$ $\leq\mathbb{E}\left(\epsilon^{\frac{p^{\star}}{2}}\sup_{t\geq 0}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}^{p^{\star}}\right)/\delta^{p^{\star}}$
$\leq\frac{\epsilon^{\frac{p^{\star}}{2}}C_{\alpha,p^{\star}}(1+\bar{R}_{r})^{p^{\star}}}{\delta^{p^{\star}}}.$
Then, taking any $\delta,\eta>0$ and $t\geq 0$ there exists $\epsilon_{r}>0$ such that for any $\epsilon<\epsilon_{r}$,
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\Gamma^{\alpha}_{\bar{R}_{r}}(t)\|_{L^{p^{\star}}}\geq\delta\right)<\eta.$ (4.12)
Here, $\epsilon_{r}$ does not depend on $t$. By Theorem 2.10 and Theorem 2.13 of [salins2019equivalences], (4.12) implies that the family of measures
$\upsilon^{\epsilon}(B):=\mathbb{P}\left(\epsilon^{\frac{1}{2}}\Gamma^{\bar{R}_{r},\alpha}_{t}(\cdot)\in B\right),\;B\in\mathcal{B}\left(C([0,t];L^{p^{\star}})\right),$
satisfies LDP in $L^{p^{\star}}$ uniformly for $t\geq 0$ with the rate functional
$I^{\bar{R}_{r},\alpha}_{t}(z)=\frac{1}{2}\inf\left\{\int_{0}^{t}\|f(s)\|_{H}^{2}\;dt;z=\Gamma^{\bar{R}_{r},\alpha}_{t}(f)\right\},$
where
$\Gamma^{\bar{R}_{r},\alpha}_{t}(f)(\tau):=\int_{0}^{\tau}(\tau-s)^{-\alpha}S(\tau-s)e^{-\lambda(\tau-s)}(1+\bar{R}_{r})f(s)\;ds.$
Denote
$K^{\bar{R}_{r},\alpha}_{t}(r):=\{z\in C([0,t];E):I^{\bar{R}_{r},\alpha}_{t}(z)\leq r\}$
to be a level set of $I^{\bar{R}_{r},\alpha}_{t}$. By the same procedure as above, for $f\in L^{2}(0,t;H)$ and $t\geq 0$, $\|\Gamma^{\bar{R}_{r},\alpha}_{t}(f)\|_{C([0,t];L^{p^{\star}})}\leq C\|f\|_{L^{2}(0,t;H)}$. This implies that for any $r>0$, $K^{\bar{R}_{r},\alpha}_{t}(r)$ is bounded in $C([0,t];L^{p^{\star}})$. Then, by the same argument as Theorem 3.9 in [BAI2026111284] and (4.10), there exist $\delta^{\star}>0$ and $\epsilon_{r}$, $\tilde{R}_{r}$ depend on $\bar{R}_{r}$, $r$ such that for any $t\geq 0$
$\ \ \ \ \mathbb{P}\left(\bigcup_{\|v\|_{C([0,t];E)}\leq\bar{R}_{r}}\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(v)(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\Gamma^{\bar{R}_{r},\alpha}_{t}(s)\|_{L^{p^{\star}}}\geq\tilde{R}_{r}\right)$
$\leq\mathbb{P}\left(\epsilon^{\frac{1}{2}}\Gamma^{\bar{R}_{r},\alpha}_{t}(\cdot)\in B_{C([0,t];L^{p^{\star}})}^{c}(K^{\bar{R}_{r},\alpha}_{t}(r),\delta)\right)$
$\leq\exp(-\frac{r}{\epsilon}),\;\epsilon\leq\epsilon_{r}.$
Hence, for $\epsilon\leq\epsilon_{r}$ and all $t\geq 0$
$\mathbb{P}\left(\epsilon^{\frac{1}{2}}\sup_{s\in[0,t]}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p^{\star}}}\geq\tilde{R}_{r},\;\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(s)\|_{E}\leq\bar{R}_{r}\right)\leq\exp(-\frac{r}{\epsilon}).$ (4.13)
Next, we need to estimate
$\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\mathbb{P}\left(\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(s)\|_{E}\geq\bar{R}_{r}\right)\;dt.$
Notice that by (3.12), the $C([0,t];E)$ norm of $z_{f}^{0}$ is bounded by $\|f\|_{L^{2}(0,t;H)}$ for all $t\geq 0$. Then, for any $r>0$, the level set $K^{0}_{t}$ is bounded in $C([0,t];E)$ uniformly with respect to all $t\geq 0$. Therefore, suppose we can prove that for any $r>0$ and all $t\geq 0$, there exists $\epsilon_{r}>0$ that does not depend on $t$, such that for any $\epsilon<\epsilon_{r}$
$\mathbb{P}\left(\text{dist}_{C([0,t];E)}(\bar{u}_{\epsilon}^{0},K^{0}_{t}(2r))\geq M\right)\leq\exp({-\frac{r}{\epsilon}})$
holds, then the proof is complete. Here, $M>0$ depends on $r$ and will be chosen later. To show that, we define functions from $C([0,t];E)$ to $\mathbb{R}$,
$h_{r,t,M}(\varphi):=2r-2r\cdot\min\left\{1,\frac{\text{dist}_{C([0,t];E)}(\varphi,K^{0}_{t}(2r))}{M}\right\},\;\varphi\in C([0,t];E).$ (4.14)
By this definition, it is clear that $h_{r,t,M}$ is bounded and Lipschitz continuous, with a Lipschitz constant of $\frac{2r}{M}$. Moreover, by the measurability of $\bar{u}^{0}_{\epsilon}$ on $C([0,t];E)$, the real valued random variable $h_{r,t,M}(\bar{u}^{0}_{\epsilon})$ is also measurable. Note that if $\text{dist}_{C([0,t];E)}(\bar{u}^{0}_{\epsilon},K^{0}_{t}(2r))\geq M$, then $h_{r,t,M}(\bar{u}^{0}_{\epsilon})=0$. Thus,
$\mathbb{P}\left(\text{dist}_{C([0,t];E)}(\bar{u}_{\epsilon}^{0},K^{0}_{t}(2r))\geq M\right)\leq\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right).$ (4.15)
Define by $\bar{u}^{0}_{\epsilon,f}$ the dynamics of the following equation on $E$
[…]
Then, by the variational representations in Theorem 3 of [budhiraja2008large], for any $\epsilon>0$
$\begin{split}\epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)=-\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}.\end{split}$ (4.17)
By the definition of infimum, for any $0<\eta\leq\min\{1,r/2\}$, there exists $f_{n}\in L^{2}(0,t;H)$ such that
$\begin{split}&\ \ \ \ \mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f_{n}})\right\}\\ &\leq\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}+\eta.\end{split}$ (4.18)
We claim that such $f_{n}$ must satisfy $\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds\leq 2r+1$. This is due to the fact that $h_{r,t,M}$ is bounded by $2r$. Indeed, if $\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds>2r+1$, then by (4.18), one gets
$\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}>2r.$
However, by taking $f=0$, one gets
$\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}\leq 2r,$
which contradicts.
Similarly, by the definition of the rate function
$\begin{split}\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}&=\inf_{f\in L^{2}(0,t;H)}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(z^{0}_{f})\right\}\\ &\leq\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(z^{0}_{f_{n}}).\end{split}$ (4.19)
By substituting (4.18) and (4.19) into (4.17), we get
$\ \ \ \ \epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)+\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}$
$\leq-\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f_{n}})\right\}+\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(z^{0}_{f_{n}})+\eta$
$\leq\mathbb{E}\left[h_{r,t,M}(z^{0}_{f_{n}})-h_{r,t,M}(\bar{u}^{0}_{\epsilon,f_{n}})\right]+\eta$
$\leq\frac{2r}{M}\mathbb{E}\|z^{0}_{f_{n}}-\bar{u}^{0}_{\epsilon,f_{n}}\|_{C([0,t];E)}+\eta.$
Here, the last inequality is due to the Lipschitz continuity of $h_{r,t,M}$. Since $\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds\leq 2r+1$, as in the proof of Proposition 6.1 of [cerrai2003stochastic], we have that there exists $C_{r}>0$ that depends on $r$ such that for any $0<\epsilon\leq 1$,
$\sup_{t\geq 0}\|z^{0}_{f_{n}}(t)\|_{E}\leq C_{r},\;\mathbb{E}\left[\sup_{t\geq 0}\|\bar{u}^{0}_{\epsilon,f_{n}}(t)\|\right]\leq C_{r}.$
Hence, by choosing $M\geq 8C_{r}$, one gets that for any $0<\epsilon\leq 1$ and all $t\geq 0$
$\epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)+\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}\leq\frac{4C_{r}}{M}\cdot r+\eta\leq r.$ (4.20)
Note that for any $\varphi\in C([0,t];E)$, either $\varphi\in K^{0}_{t}(2r)$ in which case $h_{r,t,M}(\varphi)=2r+1$, or $\varphi\notin K^{0}_{t}(2r)$ in which case $I_{t}^{0}(\varphi)>2r$. This implies that
$\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}\geq 2r.$ (4.21)
Now, substituting (4.20), (4.21) and (4.17) into (4.15) , we get for any $0<\epsilon\leq 1$ and all $t\geq 0$
$\ \ \ \ \epsilon\log\mathbb{P}\left(\text{dist}_{C([0,t];E)}(\bar{u}_{\epsilon}^{0},K^{0}_{t}(2r))\geq M\right)$
$\leq\epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)+\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}-2r$
$\leq-r,$
which is the desired result.
The proof of Proposition 4.2 is complete. ∎
4.2 The LDP lower bound
Proposition 4.4.
The family of invariant measures $\{\bar{\mu}^{\epsilon}\}_{\epsilon>0}$ of equation (1.3) satisfies LDP lower bound in $E$ with rate functional $U_{L}(\cdot)$. That is, for any $\zeta\in E,\;\delta>0$ and $\gamma>0$, there exists $\epsilon_{0}>0$ such that
$\bar{\mu}_{\epsilon}(B_{E}(\zeta,\delta))\geq\exp{\left(-\frac{U(\zeta)+\gamma}{\epsilon}\right)},\;\epsilon<\epsilon_{0}.$
Proof.
Fix $\zeta\in E$, $\delta>0$, $\gamma>0$, and $T>0$. We assume that $U(\zeta)<\infty$, or otherwise the result would be trivial. We set $\left\{z^{x}\right\}_{x\in E}\subset C([0,T];E)$ to be a family of paths satisfying
$\sup_{x\in E}\left\|z^{x}(T)-\zeta\right\|_{E}<\delta/2.$ (4.22)
By Proposition 4.2, there exists a constant $R_{r}$ being sufficiently large and independent of $\epsilon$ such that
$\bar{\mu}_{\epsilon}\left(\overline{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}\right)\geq\frac{1}{2},\ \text{for all}\ \epsilon_{0}>\epsilon>0.$ (4.23)
Moreover, according to Section 2, $\left\{\mathcal{L}(\bar{u}_{\epsilon}^{x})\right\}_{\epsilon>0}$ satisfies the Freidlin-Wentzell uniform large deviations principle in $C([0,T];E)$. So, for every $r_{0},\gamma>0$ there exists $\epsilon_{0}>0$ such that $\gamma/2+\epsilon_{0}\ln\frac{1}{2}>0$ and for any $z^{x}\in C([0,T];E)$ with $I_{T}^{x}(z^{x})\leq r_{0}$,
$\inf_{x\in B_{E}(0,R)}\mathbb{P}\left(\left\|\bar{u}_{\epsilon}^{x}-z^{x}\right\|_{C([0,T];E)}<\delta/2\right)\geq\inf_{x\in B_{E}(0,R)}\exp\left(-\frac{1}{\epsilon}\left[I_{T}^{x}(z^{x})+\gamma/2+\epsilon\ln\frac{1}{2}\right]\right),$ (4.24)
for every $\epsilon\leq\epsilon_{0}$. Since the family $\left\{\bar{\mu}_{\epsilon}\right\}_{\epsilon>0}$ is the invariant measure of equation (1.3), from (4.22), (4.23) and (4.24), we have
$\bar{\mu}_{\epsilon}\left(B_{E}(\zeta,\delta)\right)$ $=\int_{E}\mathbb{P}\left(\|\bar{u}_{\epsilon}^{x}(T)-\zeta\|_{E}<\delta\right)\,d\bar{\mu}_{\epsilon}(x)$
$\geq\int_{E}\mathbb{P}\left(\left\|\bar{u}_{\epsilon}^{x}-z^{x}\right\|_{C([0,T];E)}<\delta/2\right)\,d\bar{\mu}_{\epsilon}(x)$
$\geq\int_{\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\mathbb{P}\left(\left\|\bar{u}_{\epsilon}^{x}-z^{x}\right\|_{C([0,T];E)}<\delta/2\right)\,d\bar{\mu}_{\epsilon}(x)$
$\geq\bar{\mu}_{\epsilon}\left(\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}\right)\inf_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\mathbb{P}\left(\left\|\bar{u}_{\epsilon}^{x}-z^{x}\right\|_{C([0,T];E)}<\delta/2\right)$
$\geq\exp\left(\frac{\epsilon\ln\frac{1}{2}}{\epsilon}\right)\inf_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\exp\left(-\frac{1}{\epsilon}\left[I_{T}^{x}(z^{x})+\gamma/2+\epsilon\ln\frac{1}{2}\right]\right)$
$=\inf_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\exp\left(-\frac{1}{\epsilon}\left[I_{T}^{x}(z^{x})+\gamma/2\right]\right),$
for every $\epsilon\leq\epsilon_{0}$. Therefore, to complete the proof, it suffices to find a sufficiently large $T$ such that for each $x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}$, there exists a path $z^{x}\in C([0,T];H)$ with $z^{x}(0)=x$ that satisfies
$I_{T}(z^{x})\leq U(\zeta)+\gamma/2$
and
$\|z^{x}(T)-\zeta\|_{H}<\delta/2.$
For this, denote the solution of equation (2.3) of the skeleton function by $z^{x}_{f}$. By the definition of $U(\zeta)$, there exist $T_{2}>0$, $z_{\bar{f}}^{m_{L}-\psi}\in C([0,T];E)$ and $\bar{f}\in L^{2}([0,T];H)$ such that $z^{m_{L}-\psi}_{\bar{f}}(0)=m_{L}-\psi$, $z^{m_{L}-\psi}_{\bar{f}}(T_{2})=\zeta$ and
$\frac{1}{2}\int_{0}^{T_{2}}\|\bar{f}(t)\|_{H}^{2}\,dt=I_{T_{2}}(z^{m_{L}-\psi}_{\bar{f}})\leq U(\zeta)+\gamma/2.$
Meanwhile, by Lemma 3.3 when $f=0$, we have that for any $x\in B_{E}(r)$,
$\lim_{t\rightarrow+\infty}\|z^{x}_{0}(t)-(m_{L}-\psi)\|_{E}\leq C\lim_{t\rightarrow\infty}\|z^{x}_{0}(t)-(m_{L}-\psi)\|_{H^{1}}=0.$
Thus, by the compactness of $\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}$ in $E$ and (2.8), for any $\eta>0$ we can take $T_{1}=T_{1}(\eta)$ large enough such that
$\sup_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\left\|z^{x}_{0}(T_{1})-(m_{L}-\psi)\right\|_{E}\leq\eta.$
We choose the paths $z^{x}=z^{x}_{f}$ with $f\in L^{2}([0,T];H)$ defined by
$f(t):=\begin{cases}0,&t\in[0,T_{1}),\\ \bar{f}(t-T_{1}),&t\in[T_{1},T_{1}+T_{2}],\end{cases}$
and $T=T_{1}+T_{2}$. Clearly, we have
$I_{T}(z^{x})$ $=\frac{1}{2}\int_{0}^{T}\|f(t)\|_{H}^{2}\,dt=\frac{1}{2}\int_{0}^{T_{2}}\|\bar{f}(t)\|_{H}^{2}\,dt\leq U(\zeta)+\gamma/2.$
Moreover, from the above construction and (2.8) we have that
$\sup_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\left\|z^{x}_{f}(T)-\zeta\right\|_{E}$ $\leq\sup_{y\in B_{E}(m_{L}-\psi,\lambda)}\left\|z^{y}_{\bar{f}}(T_{2})-z^{m_{L}-\psi}_{\bar{f}}(T_{2})\right\|_{E}$
$\leq\sup_{y\in B_{E}(m_{L}-\psi,\lambda)}c(T,\|y\|_{E},U(\zeta))\cdot\|y\|_{E}.$
Therefore, we can choose $\lambda$ small enough such that
$\|z^{x}_{f}(T)-\zeta\|_{E}<\delta/2.$
In conclusion, the path $z^{x}(t)=z^{x}_{f}(t)$ satisfies the two desired conditions. The proof is complete.
4.3 The LDP upper bound
To prove the upper bound for LDP, we need the following two auxiliary results.
Lemma 4.5.
Under Assumption 1, for any $\delta>0$ and $r>0$, there exist $\beta>0$ and $T>0$ such that for any $t^{\star}\geq T$ and $z\in C([-t^{\star},0];E)$,
$\|z(-t^{\star})-(m_{L}-\psi)\|_{E}<\beta,\ I_{t^{\star}}(z)\leq r\Rightarrow dist_{E}\left(z(0),\Phi_{L}(r)\right)<\delta,$
where
$\Phi_{L}(r):=\left\{x\in E:U_{L}(x)\leq r\right\}.$
Proof.
Assume the claim does not hold. Then, there exist $\delta>0$, $r>0$ and a sequence of functions $z_{n}\in C([-T_{n},0];E)$ with $T_{n}\nearrow+\infty$ such that $\beta_{n}:=\|z(-T_{n})-(m_{L}-\psi)\|_{E}\searrow 0$ and
$I_{T_{n}}(z_{n})\leq r\ ,dist_{H}(z_{n}(0),\Phi_{L}(r))\geq\delta.$ (4.25)
Thus, $z_{n}$ also satisfies equation (2.3) with $z_{n}(-T_{n})=x$ and $f$ such that $\|f\|_{L^{2}([-T_{n},0];H)}^{2}=2I_{T_{n}}(z_{n})\leq 2r$.
The following arguments are similar to the proof of Proposition 3.6. By Lemma 3.5, for any $k\in[0,T_{n}]$ and $k\in\mathbb{N}$, the restriction of $z_{n}$ to the interval [k,0][-k,0] belongs to $K_{\Lambda,-k,0}$, where
$\Lambda:=\{\bar{u}\in E;\mathbf{E}^{*}_{L}(\bar{u})\leq C_{r}\},$
and $C_{r}$ is a constant depending on $r$. As the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$, we find that $K_{\Lambda,-k,0}$ are compact in C([k,0];E)C([-k,0];E). Taking $k=1$, we can find $\{z_{n_{1}}\}\subseteq\{z_{n}\}$ and $\hat{z}_{1}\in C([-1,0];E)$ such that $z_{n_{1}|_{[-1,0]}}$ converges to $\hat{z}_{1}$ in C([1,0];E)C([-1,0];E). With the same arguments, we can find a subsequence $\{z_{n_{2}}\}\subset\{z_{n_{1}}\}$ and $\hat{z}_{2}\in C([-2,0];E)$ such that $z_{n_{2}|_{[-2,0]}}$ converges to $\hat{z}_{2}$ in C([2,0];E)C([-2,0];E). Continuing this procedure, we can find a subsequence $\{z_{n^{\prime}}\}\subset\{z_{n}\}$ converging to some $\hat{z}$ in $C((-\infty,0];E)$. Thanks to Theorem 5.1 of [cerrai2004large], $I_{-k}(\cdot)$ is lower semi-continuous, so we have $I_{-k}(\hat{z})\leq r$ for any $k\geq 0$ which implies $I_{-\infty}(\hat{z})\leq r$. Moreover, as $\beta_{n}\searrow 0$, we have
$\lim_{t\rightarrow+\infty}\|\hat{z}(-t)-(m_{L}-\psi)\|_{E}=0.$
Therefore, $\hat{z}\in K_{-\infty}(r)$. Thus, by the characterisation of $U_{L}$ (see Proposition 3.8), $U_{L}(\hat{z}(0))\leq r$ and $\hat{z}(0)\in\Phi_{L}(r)$. However, by our construction
$\lim_{n_{1}\rightarrow\infty}\|z_{n_{1}}(0)-\hat{z}(0)\|_{E}=0,$
which contradicts (4.25). The proof of the desired result is complete.
Lemma 4.6.
For any $r>0$, $\delta>0$ and $\rho>0$, let $\beta(r,\delta)$ be as in Lemma 4.5. Then there exists $\bar{N}\in\mathbb{N}$ large enough such that
$u\in H_{\rho,r,\delta}(\bar{N})\ \text{implies}\ I_{\bar{N}}(u)\geq r,$
where the set $H_{\rho,r,\delta}(\bar{N})$ is defined for $\bar{N}\in\mathbb{N}$ by
$H_{\rho,r,\delta}(\bar{N}):=\left\{u\in C([0,\bar{N}];E),\|u(0)\|_{E}\leq\rho,\|u(j)-(m_{L}-\psi)\|_{E}\geq\beta,j=1,...,\bar{N}\right\}.$
Proof.
Assume there exist $r>0$, $\delta>0$, and $\rho>0$ such that for any $n\in\mathbb{N}$ there exists $u\in H_{\rho,r,\delta}(n)$ with
$I_{n}(u)\leq r.$ (4.26)
Then, by Proposition 3.2 and (3.12), $\mathbf{E}^{*}_{L}(u(n))\leq C_{r,\rho}$, for all $n\in\mathbb{N}$. Here, $C_{r,\rho}$ is a positive constant depending on $r$ and $\rho$. Now, for any $k\in\mathbb{N}$ we define
$r_{k}:=\inf\left\{I_{k}(u):u\in C([0,k];E),\mathbf{E}^{*}_{L}(u(0))\leq C_{r,\rho},\|u(k)-(m_{L}-\psi)\|_{E}\geq\beta\right\}.$
Notice that if $u\in H_{\rho,r,\delta}(n)$, then for any $k\in\mathbb{N}$, $k\leq n$, $\mathbf{E}^{*}_{L}(u(k))\leq C_{r,\rho}$ and $\|u(k)-(m_{L}-\psi)\|_{E}\geq\beta$. Therefore, if we can show that there exists $\bar{k}\in\mathbb{N}$ such that $r_{\bar{k}}>0$, then for any $u\in C([0,n\bar{k}];E)$
$I_{n\bar{k}}(u)\geq nr_{\bar{k}},$
which contradicts (4.26). To finish our proof, it remains to show that there exists $\bar{k}\in\mathbb{N}$ such that $r_{\bar{k}}>0$. In fact, if $r_{\bar{k}}=0$ for all $\bar{k}$, then there exists a sequence
$\{u_{i}\in C([0,\bar{k}];E),\mathbf{E}^{*}_{L}(u_{i}(0))\leq C_{r,\rho},\|u_{i}(\bar{k})-(m_{L}-\psi)\|_{E}\geq\beta\},i\in\mathbb{N},$
such that
$\lim_{i\rightarrow\infty}I_{\bar{k}}(u_{i})=0.$ (4.27)
By the same arguments as in the proof of Lemma 3.7, for any $\eta>0$, there exist $\bar{k}$ sufficiently large, $N\in\mathbb{N}$ and some $t\in[\frac{\bar{k}}{2},\bar{k}]$ such that for any $i\in\mathbb{N}$, $i\geq N$
$\mathbf{E}^{*}_{L}(u_{i}(t))\leq\frac{\eta}{2}.$
Furthermore, by (3.11) and (3.23),
$\mathbf{E}^{*}_{L}(u_{i}(\bar{k}))$ $\leq\mathbf{E}^{*}_{L}(u_{i}(t))+C(1+\sup_{s\geq 0,i\in\mathbb{N}}\|u_{i}(s)\|_{E}^{2})\cdot I_{\bar{k}}(u_{i})$
$\leq\mathbf{E}^{*}_{L}(u_{i}(t))+CI_{\bar{k}}(u_{i}).$
Since $\lim_{i\rightarrow\infty}I_{\bar{k}}(u_{i})=0$, for any $\eta>0$ there exists $N$ such that for all $i\geq N$, $CI_{\bar{k}}(u_{i})\leq\frac{\eta}{2}$. Therefore, one obtains that, for $i$ sufficiently large,
$\mathbf{E}^{*}_{L}(u_{i}(\bar{k}))\leq\eta.$
Moreover, by Proposition 3.1, for any $\beta>0$, one can choose $\eta$ sufficiently small such that
$\|u_{i}(\bar{k})-(m_{L}-\psi)\|_{E}\leq\frac{\beta}{2}$
which contradicts the assumption that $\|u_{i}(\bar{k})-(m_{L}-\psi)\|_{E}\geq\beta$. The proof is complete. ∎
Theorem 4.7.
Under Assumption 1, the family of invariant measures $\left\{\bar{\mu}_{\epsilon}\right\}_{\epsilon>0}$ satisfies a large deviation principle upper bound in $E$ with the rate function $U_{L}$. That is, for any $r\geq 0$ and $\delta,\gamma>0$, there exists $\epsilon_{0}>0$ such that
$\bar{\mu}_{\epsilon}\left(\left\{h\in E:dist_{E}(h,\Phi_{L}(r))\geq\delta\right\}\right)\leq\exp\left(-\frac{r-\gamma}{\epsilon}\right),\ \epsilon<\epsilon_{0}.$
Proof.
For any fixed $r>0$, $\delta>0$ and $\gamma>0$, let $R_{r}$, $\epsilon_{r}$ $k^{\star}$ and $p^{\star}$ be as in Proposition 4.2. Since $\bar{\mu}_{\epsilon}$ is the invariant measure for equation (1.3), for $t\geq\bar{N}$ we have
$\ \ \ \ \ \bar{\mu}_{\epsilon}\left(\left\{h\in E:dist_{E}(h,\Phi_{L}(r))\geq\delta\right\}\right)$
$=\int_{E}\mathbb{P}\left(dist_{E}(\bar{u}_{\epsilon}^{y}(t),\Phi_{L}(r))\geq\delta\right)\,d\bar{\mu}_{\epsilon}(y)$
$=\int_{B_{W^{k^{\star},p^{\star}}}^{c}(0,R_{r})}\mathbb{P}\left(dist_{E}(\bar{u}_{\epsilon}^{y}(t),\Phi_{L}(r))\geq\delta\right)\,d\bar{\mu}_{\epsilon}(y)$
$\ \ \ \ +\int_{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}\mathbb{P}\left(dist_{E}(\bar{u}_{\epsilon}^{y}(t),\Phi_{L}(r))\geq\delta,u_{\epsilon}^{y}\in H_{R_{r},r,\delta}(\bar{N})\right)\,d\bar{\mu}_{\epsilon}(y)$
$\ \ \ \ +\int_{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}\mathbb{P}\left(dist_{E}(\bar{u}_{\epsilon}^{y}(t),\Phi_{L}(r))\geq\delta,\bar{u}_{\epsilon}^{y}\notin H_{R_{r},r,\delta}(\bar{N})\right)\,d\bar{\mu}_{\epsilon}(y)$
$=:K_{1}+K_{2}+K_{3},$
where $\bar{N}$ is chosen as in Lemma 4.6 such that
$u\in H_{R_{r},r,\delta}(\bar{N})\ \text{implies}\ I_{\bar{N}}(u)\geq r.$
Now, thanks to Proposition 4.2 we have
$K_{1}\leq\bar{\mu}_{\epsilon}\left(B^{c}_{W^{k^{\star},p^{\star}}}(0,R_{r})\right)\leq\exp(-\frac{r}{\epsilon}),\ \epsilon\leq\epsilon_{r}.$
To estimate $K_{2}$, notice that $\overline{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}$ is a compact subset of E, and $H_{R_{r},r,\delta}(N)$ is a closed set in $C([0,N];E)$. Then, by Corollary 2.2, $\left\{\mathcal{L}(\bar{u}_{\epsilon}^{x})\right\}_{\epsilon>0}$ satisfies a Dembo-Zeitouni uniform large deviation principle over compact sets, we infer that there exists $\epsilon_{1}>0$ such that
$K_{2}$ $\leq\sup_{y\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}}\mathbb{P}\left(\bar{u}_{\epsilon}^{y}\in H_{R_{r},r,\delta}(\bar{N})\right)$
$\leq\exp\left(-\frac{1}{\epsilon}\left[\inf_{z\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}}\inf_{h\in H_{R_{r},r,\delta}(\bar{N})}I_{N}^{z}(h)-\gamma\right]\right)$
$\leq\exp\left(-\frac{r-\gamma}{\epsilon}\right),$
for any $\epsilon\leq\epsilon_{1}$. Here, the third inequality is due to Lemma 4.6.
Finally, let us deal with the third term $K_{3}$. By the definition of $H_{R_{r},r,\delta}(N)$ and the Markov property of $u_{\epsilon}$, we have
$K_{3}$ $=\int_{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}\mathbb{P}\left(\bigcup_{j=1}^{\bar{N}}\left\{\|\bar{u}_{\epsilon}^{y}(j)\|_{E}<\beta\right\}\bigcap\left\{dist_{E}(\bar{u}_{\epsilon}^{y}(t),\Phi_{L}(r))\geq\delta\right\}\right)\,d\bar{\mu}_{\epsilon}(y)$
$\leq\sum_{j=1}^{N}\int_{B_{W^{k^{\star},p^{\star}}}(0,R_{r})}\mathbb{P}\left(\left\{\|\bar{u}_{\epsilon}^{y}(j)\|_{E}<\beta\right\}\bigcap\left\{dist_{E}(\bar{u}_{\epsilon}^{y}(t),\Phi_{L}(r))\geq\delta\right\}\right)\,d\bar{\mu}_{\epsilon}(y)$
$\leq\sum_{j=1}^{N}\sup_{y\in B_{E}(0,\beta)}\mathbb{P}\left(dist_{E}(\bar{u}_{\epsilon}^{y}(t-j),\Phi_{L}(r))\geq\delta\right).$
In order to use the ULDP, we need to transfer the event at tjt-j to an event in C([0,tj];E)C([0,t-j];E). To achieve that, we choose $t\geq T+N$, where $T$ was given in Lemma 4.5. We then have that for $y\in B_{E}(0,\lambda)$, if
$dist_{C([0,t-j];E)}\left(\bar{u}_{\epsilon}^{y},K^{y}(r)\right)<\frac{\delta}{2},$
then
$\inf\left\{\|\bar{u}_{\epsilon}^{y}-v\|_{C([0,t-j];E)}:v\in C([0,t-j];E),\ \|v(0)\|_{E}<\lambda,\ I_{t-j}(v)\leq r\right\}<\frac{\delta}{2},$
where
$K^{y}(r):=\left\{v\in C([0,t-j];E):v(0)=y,I_{t-j}(v)\leq r\right\}.$
It turns out by applying Lemma 4.5 for $\frac{\delta}{2}$ that
$dist_{E}(\bar{u}_{\epsilon}^{y}(t-j),\Phi_{L}(r))<\delta.$
Thus, according the FWULDP results, there exists $\epsilon_{0,j}>0$ such that for any $\epsilon\leq\epsilon_{0,j}$,
$\sup_{y\in B_{E}(0,\lambda)}\mathbb{P}\left(dist_{E}(\bar{u}_{\epsilon}^{y}(t-j),\Phi_{L}(r))\geq\delta\right)$ $\leq\sup_{y\in B_{E}(0,\lambda)}\mathbb{P}\left(dist_{C([0,t-j];E)}\left(\bar{u}_{\epsilon}^{y},K^{y}(r)\right)\geq\frac{\delta}{2}\right)$
$\leq\exp\left(-\frac{r-\gamma}{\epsilon}\right).$
Therefore, we choose $\epsilon_{0}=\min(\epsilon_{r},\epsilon_{1},\epsilon_{0,1},...,\epsilon_{0,N})>0$, it follows that for any $\epsilon\leq\epsilon_{0}$
$K_{3}\leq\bar{N}\exp\left(-\frac{r-\gamma}{\epsilon}\right),$
Combining the estimates of $K_{1}$, $K_{2}$, $K_{3}$, we complete the proof of this proposition. ∎
With the upper bound of LDP for $\bar{\mu}_{\epsilon}$, it is easy to see the following result.
Corollary 4.8.
Under Assumption 1, there exist $C>0$ and $\delta_{0}>0$ such that for every $0<\delta<\delta_{0}$ there exists $\epsilon_{0}>0$ such that
$\mu_{\epsilon}(\{h\in\mathcal{E}:\|h-m_{L}\|_{E}\geq\delta\})\leq\exp\left(-\frac{C\delta^{2}}{\epsilon}\right),\epsilon<\epsilon_{0}.$
In particular, the measures $\mu_{\epsilon}$ concentrate around the minimiser $m_{L}$ exponentially fast, when $\epsilon$ is sufficiently small.
Proof.
This is a direct consequence of Propositions 3.1 and 3.4, and the LDP upper bound for $\bar{\mu}_{\epsilon}$ given in Proposition 4.7. ∎
Appendix Acknowledgements
[…]